### v1

# Efficient Action Counting with Dynamic Queries

Abstract
Temporal repetition counting aims to quantify the repeated action cycles within a video. The majority of existing methods rely on the similarity correlation matrix to characterize the repetitiveness of actions, but their scalability is hindered due to the quadratic computational complexity. In this work, we introduce a novel approach that employs an action query representation to localize repeated action cycles with linear computational complexity. Based on this representation, we further develop two key components to tackle the essential challenges of temporal repetition counting. Firstly, to facilitate open-set action counting, we propose the dynamic update scheme on action queries. Unlike static action queries, this approach dynamically embeds video features into action queries, offering a more flexible and generalizable representation. Secondly, to distinguish between actions of interest and background noise actions, we incorporate inter-query contrastive learning to regularize the video representations corresponding to different action queries. As a result, our method significantly outperforms previous works, particularly in terms of long video sequences, unseen actions, and actions at various speeds. On the challenging RepCountA benchmark, we outperform the state-of-the-art method TransRAC by 26.5% in OBO accuracy, with a 22.7% mean error decrease and 94.1% computational burden reduction. Code is available at https://github.com/lizishi/DeTRC.
Keywords:
Temporal repetition counting Video understanding
1 Introduction
Temporal periodicity is a ubiquitous phenomenon in the natural world. Temporal Repetition Counting (TRC) aims to accurately measure the number of repetitive action cycles within a given video and holds significant potential for applications ranging from pedestrian detection [29] to fitness monitoring [11].
Pioneer methods  [16, 1, 7, 38, 27, 37, 6] represent time-series video data as one-dimensional signals and employ spectral analysis techniques such as the Fourier transform. While suitable for short videos with fixed periodic cycle lengths, these methods struggle to handle real-world scenarios with varying cycle lengths and sudden interruptions. Recent studies shift to deep learning-based methods [17, 9, 13, 46] and show promising performances. Notably, most of these methods, such as RepNet [9] and TransRAC [13], utilize a temporal similarity correlation matrix to depict repetitiveness, as illustrated in Fig. 1 (a). Nevertheless, the computational complexity of this representation grows quadratically with the number of input frames $T$, highlighting a significant gap in scalability that hinders their application to real-world scenarios of varying action periods and dynamics.
Recent progress in action detection [33, 22] introduces an efficient representation of action periods by associating each action instance with an action query, similar to DETR [3]. Inspired by this, we propose to formulate the TRC problem as a set prediction task where the goal is to detect every action cycle by representing it as an action query. This novel formulation reduces the complexity from quadratic to linear and enables counting long videos with varying action periods, as illustrated in Fig. 1 (b). However, directly applying the action detection approaches [22, 44, 33] to the TRC problem proves inadequate (Tab. 1) in addressing two distinctive challenges unique to TRC. These challenges underscore the complexity of TRC, highlighting why TRC is not merely another action detection task but requires a nuanced approach that considers the unique nature of repetitive actions. We highlight the two inherent differences between TRC and the classical action detection task:
  1. 1.
    TRC requires recognizing open-set action instances depending on the input video, rather than detecting predefined action classes in the detection task.
  2. 2.
    TRC requires recognizing action instances with identical content, while detection does not.
Figure 1: Comparison of (a) similarity-matrix-based methods and (b) our proposed action query-based method. Unlike the similarity-matrix-based methods, which exhibit a quadratic growth $O(T^{2}C)$ in computational complexity as the length of the input video $T$ increases, our approach is characterized by a linear growth $O(TC)$ in complexity, where $C$ denotes the feature dimension of the video frames. This distinct advantage allows us to effectively process long videos that contain varying durations of action sequences.
As a result, approaching TRC as a simple action detection task results in inferior performance as shown in Sec. 4.3. In contrast, we propose two novel strategies to address these challenges and redefine the framework for TRC. In response to the first challenge, we propose Dynamic Action Query (DAQ), which adaptively updates the action query using distilled content features from the video encoder. This mechanism allows the decoder to attend to the “action of interest” based on the input video contents in a dynamic and contextually aware manner, without the need for manual definition. To tackle the second challenge, we propose Inter-query Contrastive Learning (ICL). It enforces that the repetitive action cycles are grouped together in the learned representation space while being separated apart from non-repetitive (e.g. background or distracting) video content. The integration of two core components (DAQ and ICL) ensures that the action instances are identified adaptively based on the video content and their contextual similarity. In other words, our queries are designed to localize the contextually similar action instances, which aligns exactly with the definition of repetition counting. Extensive experiments validate the effectiveness of the two proposed designs.
We summarize our contributions as follows:
  1. 1.
    We provide a novel perspective to tackle the TRC problem using a simple yet effective representation for action cycles. Our approach reduces the computational complexity from quadratic to linear and is robust to varying action periods and video lengths.
  2. 2.
    We propose Dynamic Action Query to guide the model to focus on the action of interest and improve generalization ability across different actions.
  3. 3.
    We introduce Inter-query Contrastive Learning to facilitate learning repetitive action representations and to distinguish these from distractions.
  4. 4.
    By addressing the scalability issue and unique challenges of TRC, our method on two challenging benchmarks notably surpasses state-of-the-art (SOTA) methods in terms of both accuracy and efficiency. Notably, our method strikes an effective balance in handling various action periods and video lengths, offering a significant leap forward in the practical application of TRC technologies.
2 Related Work
2.1 Temporal Repetition Counting
Traditional methods [16, 1, 7, 38, 27, 37, 6] frequently employ spectral or frequency domain techniques for the analysis of repetitive sequences, thereby preserving the underlying repetitive motion structures. While these conventional approaches are capable of effectively handling simple motion sequences or those characterized by fixed periodicity, they prove inadequate when confronted with non-stationary motion sequences encountered in real-world scenarios. In contrast, deep-learning-based approaches [17, 9, 13, 46] have demonstrated remarkable performance improvements. Notably, RepNet [9] and TransRAC [13] leverage temporal similarity matrices of actions to construct models for counting temporal repetitions. However, these similarity-matrix-based methods are not scalable for long videos due to their quadratic computational complexity. Another research line involves predicting the start and end points of each cycle [46] from coarse to fine. Nevertheless, its practicality is hindered by the requirement for over 30 forward passes to count iteratively from a single video. In this paper, we introduce an effective action cycle representation by leveraging a Transformer encoder-decoder, which reduces the computational complexity from quadratic to linear and demonstrates superior performance in handling both fast and slow actions.
2.2 Temporal Action Detection
The field of temporal action detection [30, 47, 5, 44, 20] is typically classified into two categories: anchor-based methods, and anchor-free methods. Anchor-based methods [43, 18, 28] generate multiple anchors, subsequently classifying these anchors to determine the action boundaries. Anchor-free methods  [2, 34, 42, 19] predict action instances by directly regressing the boundary and the center point of an action instance. With the rapid development of Transformer technology, DETR [3] is introduced for object detection task [48, 26, 21, 45] and gains increasing popularity with promising performance. This paradigm promotes the study in many fields such as the action detection tasks [22, 36, 39, 41]. These methods establish a direct connection between action queries and the predicted action instances, enabling them to accurately predict the temporal boundaries of actions. Inspired by these promising results, we explore the possibility of utilizing a novel action query to represent the action cycle in TRC task. In contrast to existing action detection methods, our approach allows the model to capture the inherent repetitive content of an action cycle without relying on predefined class labels and effectively addresses confounding factors such as non-repetitive video backgrounds. This makes our approach well-suited for tackling the challenges of the TRC problem.
3 Method
3.1 Overview
Given an RGB video sequence $\mathbf{V}$ with $T$ frames, The TRC task aims to predict an integer $N$ indicating the number of estimated repetitive action cycles. Our model consists of a video backbone network $\Phi(\cdot)$, a Transformer network with encoder $\mathcal{E}(\cdot)$, decoder $\mathcal{D}(\cdot)$, and multiple prediction heads as shown in Fig. 2. The backbone network $\Phi(\cdot)$ takes a sequence of $T$ video frames as input and extracts feature vectors $\mathbf{F}\in\mathbb{R}^{T\times C}$ for each frame, where $C$ denotes the feature dimension. In practice, we utilize the commonly-used backbone works, e.g. TSN [40], and freeze the backbone parameters.
Figure 2: Framework overview. Given a video input, we first extract video features $\mathbf{F}$ with the backbone network Φ()normal-$\Phi(\cdot)$. Then we leverage a Transformer encoder-decoder network to process the features $\mathbf{F}$ to detect each action cycle instance by predicting its action class label $\mathbf{C}^{\text{cls}}$ and its time position $\mathbf{P}$. To tackle the challenge of open-set action categories, we propose Dynamic Action Query (DAQ), allowing the method to attend to the “action of interest” based on the input video contents in a dynamic and contextually aware manner. To recognize actions with identical content, we propose Inter-query Contrastive Learning (ICL) to group the “action of interest” and be separated apart from noise actions (e.g. background or distractions). Integrating DAQ and ICL, our method can identify contextually similar action instances that are adaptively conditioned on the video content, without manual action class definition. By counting the queries classified as the “action of interest” with classification confidence score $\mathbf{C}^{\text{conf}}$ exceeding the threshold $\alpha$, we finally get the total count value $N$.
We then design a Transformer encoder-decoder architecture to process the features $\mathbf{F}$ to detect each action cycle instance and obtain the final action cycle count value $N$. This requires the model to discern each detected action cycle class $\mathbf{C}^{\text{cls}}$ and its time position $\mathbf{P}$. To achieve this, we introduce the concept of action queries within the decoder $\mathcal{D}(\cdot)$, which are comprised of content queries $\mathbf{Q}^{\text{cnt}}\in\mathbb{R}^{Q\times C}$ and position queries $\mathbf{Q}^{\text{pos}}\in\mathbb{R}^{Q\times C}$, where $Q$ represents the number of queries. Refer to Fig. 2 for illustration. The content queries $\mathbf{Q}^{\text{cnt}}$ are responsible for capturing the action category, while the position queries $\mathbf{Q}^{\text{pos}}$ handle the action interval estimation. We use two prediction heads including a classification head and a position head after the decoder to predict the class label $\mathbf{C}^{\text{cls}}$ and time position $\mathbf{P}$ for each query, respectively.
To tackle the challenge of open-set action categories, we define $K=2$ classes, namely “action of interest” and “null ($\emptyset$)”. We then propose the Dynamic Action Query (DAQ) strategy which adaptively updates the content query $\mathbf{Q}^{\text{cnt}}$ of the decoder using distilled content features that are selected from the encoder $\mathcal{E}(\cdot)$. This mechanism eliminates the need for a manual definition of the specific semantics of “action of interest” and allows the decoder to attend to the “action of interest” conditioned on the input video contents in a dynamic and contextually-aware manner. Nevertheless, videos may contain distracting factors such as background elements or other noise actions, necessitating that the detected “action of interest” instances exhibit identical or similar contents. To this end, we propose Inter-query Contrastive Learning (ICL) to cluster content queries into positive action set ($S^{+}$) and negative null set ($S^{-}$) based on their predicted class label $\mathbf{C}^{\text{cls}}$. Integrating DAQ and ICL allows our method to identify contextually similar action instances that are adaptively based on the video content, and exclude other distracting actions at the same time. Finally, by counting the queries classified as the “action of interest” with classification confidence score $\mathbf{C}^{\text{conf}}$ exceeding a threshold $\alpha$, we get the total count value $N$.
3.2 Model Architecture
Encoder. The encoder $\mathcal{E}(\cdot)$ is a classical Transformer [39] architecture which processes the video features $\mathbf{F}$ extracted by the backbone network. The output features $\mathbf{H}^{\text{enc}}$ of $\mathcal{E}(\cdot)$ will then be fed into a query selection module to select $Q$ queries to be the initial content queries $\mathbf{Q}^{\text{cnt}}$ for the decoder. For clarity, we will detail the query selection module later in this section. At now, it can be considered as a rough filter that selects top-$Q$ relevant features as the input queries for decoder $\mathcal{D}(\cdot)$.
Decoder. The decoder $\mathcal{D}(\cdot)$ is also a classical Transformer [39] architecture, which takes the encoder features $\mathbf{H}^{\text{enc}}$ and two types of action queries as input and outputs a processed feature $\mathbf{H}^{\text{dec}}\in\mathbb{R}^{Q\times C}$. The action queries are constructed by content queries $\mathbf{Q}^{\text{cnt}}\in\mathbb{R}^{Q\times C}$ and position queries $\mathbf{Q}^{\text{pos}}\in\mathbb{R}^{Q\times C}$, where $\mathbf{Q}^{\text{cnt}}$ and $\mathbf{Q}^{\text{pos}}$ are designed for representing the content features and the time position of each action cycle, respectively. The content queries $\mathbf{Q}^{\text{cnt}}$ are directly initialized from the distilled content features selected from the encoder $\mathcal{E}(\cdot)$, i.e. the proposed DAQ strategy, which will be detailed in Sec. 3.3. The position queries $\mathbf{Q}^{\text{pos}}$ are initialized as learnable parameters and transformed to the initial time position $\mathbf{P}^{\rm init}=(\mathbf{m}^{\rm init},\mathbf{d}^{\rm init})\in\mathbb{R}^{Q\times 2}$ with a linear layer before feeding into the decoder $\mathcal{D}(\cdot)$, where we propose to use the midpoint $\mathbf{m}$ and the period duration $\mathbf{d}$ to denote each action time position. For clarity, we omit the linear layer in Fig. 2.
Prediction heads. The prediction heads consist of a classification head network and a position head network, which are both MLPs.
The classification head takes the output features of the decoder $\mathbf{H}^{\text{dec}}$ as input and predicts the action class label for each content query. Given that the action category is an open set and not preset, we set the number of classes $K$ to be 2 in the classification head, indicating whether it is an “action of interest” or a “null ($\emptyset$)” class. The classification head employs the Softmax function to calculate the probabilities $\mathbf{C}^{\text{cls}}$ of $K$ categories for each action query. We then designate the category with the highest probability as the predicted class label for the respective action query. Subsequently, we regard this probability value as the estimated confidence score $\mathbf{C}^{\text{conf}}$. We define an action query instance as “action of interest” when its confidence score $\mathbf{C}^{\text{conf}}$ is higher than the threshold $\alpha$.
We apply binary Cross-entropy loss to supervise the predicted class labels $\mathbf{C}^{\text{cls}}$ for each action instance, aiming to guide the model in effectively discriminating the action instances of interest from others:
$\mathcal{L}_{\text{cls}}=\sum\limits_{q=1}^{Q}{\rm CrossEntropyLoss}\left(\mathbf{C}^{\text{cls}}_{q},\tilde{\mathbf{C}}^{\text{cls}}_{q}\right)=\sum\limits_{q=1}^{Q}\sum\limits_{i=1}^{K}-\tilde{c}_{q,i}{\rm log}\left(c_{q,i}\right),$ (1)
where $\mathbf{C}^{\text{cls}}_{q}=[c_{q,1},c_{q,2},...,c_{q,K}]\in\mathbb{R}^{K}$ represents the predicted class label for $q^{th}$ action query. $\tilde{}$ denotes the GT label.
The position head estimates the residual of the midpoint and duration based on the decoder features $\mathbf{H}^{\text{dec}}$ for each query, and adds to the initial time position $\mathbf{P}^{\rm init}$ to get the final estimated time position. We supervise the predicted time positions with $L_{1}$ loss and gIoU [31] loss:
$\begin{split}\mathcal{L}_{\text{pos}}&=\sum\limits_{q=1}^{Q}\left\|\mathbf{P}_{q}-\tilde{\mathbf{P}}_{q}\right\|+\lambda_{\text{gIoU}}\sum\limits_{q=1}^{Q}\left(1-{\rm gIoU}(\mathbf{P}_{q},\tilde{\mathbf{P}}_{q})\right),\end{split}$ (2)
where $\mathbf{P}_{q}=(\mathbf{m}_{q},\mathbf{d}_{q})$ represents the predicted action instance position for the $q^{th}$ query. $\tilde{}$ stands for GT label.
Note that the $Q$ predicted instances from the prediction heads are in a random order. Therefore, before applying the two loss terms, we have to establish a one-to-one correspondence between the disordered predicted instances and the ground-truth (GT) labels. Concretely, we employ the Hungarian matching algorithm [15] and the matching process considers two criteria: 1) the action class label is the “action of interest”, and 2) the time position of the action must align with a corresponding ground-truth time position label. A matching cost matrix is constructed based on these criteria. For the first criterion, the Negative Log-Likelihood (NLL) loss is employed to assess classification accuracy. For the second aspect, both Intersection over Union (IoU) and $L_{1}$ distance are utilized to compute costs, measuring the alignment between the estimated time positions and GT labels. The matching cost matrix is of dimensions $Q\times N_{gt}$, where $N_{gt}$ represents the number of GT action instances. Utilizing the Hungarian matching algorithm [15] based on this cost matrix, we derive the prediction-to-GT matching results.
Query selection. As we mentioned above, videos may contain multiple action categories or other distracting factors such as background elements, therefore, we design a query selection module after the encoder $\mathcal{E}(\cdot)$ to filter the “action of interest” coarsely. To achieve this, we reuse the two prediction heads to predict the class labels and time positions for the output features $\mathbf{H}^{\text{enc}}$ of the encoder $\mathcal{E}(\cdot)$. Based on their predicted confidence scores $\mathbf{C}^{\text{conf}}$, the query selection module selects top-$Q$ queries with high confidence that are classified as the “action of interest” to be the initial content queries $\mathbf{Q}^{\text{cnt}}$ for the decoder. For clarity, we omit the encoder prediction heads in Fig. 2.
Note that the two prediction head networks are shared between the encoder and decoder. In addition, we apply corresponding supervision using $\mathcal{L}_{\text{cls}}$ and $\mathcal{L}_{\text{pos}}$ to the encoder prediction heads. Additional experimental results of the encoder prediction heads can be found in the supplementary materials.
3.3 Dynamic Action Query
As discussed in Sec. 1, the TRC problem requires recognizing open-set action instances depending on the video content, where the action category is not predefined. Therefore, we propose the Dynamic Action Query strategy, which adaptively updates the content query $\mathbf{Q}^{\text{cnt}}$ using distilled content features from the encoder. Simple yet effective, the DAQ strategy directly initializes the input content queries $\mathbf{Q}^{\text{cnt}}$ of the decoder using the $Q$ content queries filtered by the query selection module, as shown in Fig. 2. This mechanism provides priors for the decoder that allows it to attend to the “action of interest” conditioned on the input video contents in a dynamic and contextually aware manner. In this way, we avoid the need for the manual definition of the action categories, thereby improving the models’ generalization capability.
Furthermore, we extensively explore different methods for initializing the queries in the decoder $\mathcal{D}(\cdot)$ in the supplementary material, confirming the DAQ strategy as the most effective.
3.4 Inter-query Contrastive Learning
Since the input video may contain multiple action categories and other distractors such as the background motion, another unique challenge to the TRC task is to recognize action instances with identical content. This requires that the actions of interest we classify exhibit similarity in their motion patterns, while other action queries should have dissimilar representations. To tackle this challenge, we propose Inter-query Contrastive Learning to distinguish the action queries. Intuitively, we hope to partition the content queries into $K=2$ main categories through contrastive learning. One category comprises the “actions of interest”, with similar representations, while the other consists of null instances. The latter may encompass interfering actions or background elements, with distinct representations from those of the actions of interest.
Concretely, we designate the $Q$ decoder features $\mathbf{H}^{\text{dec}}$ features classified into two sets according to the predicted class labels $\mathbf{C}^{\text{cls}}$, as shown in Fig. 2. The features that are classified as actions of interest form the positive set $S^{+}$, while the remaining features form the negative set $S^{-}$. Then we apply contrastive learning using InfoNCE loss [12] over the feature space:
$\begin{split}\mathcal{L}_{+}&=\sum\limits_{s\in S^{+},s\neq q}{\rm exp}(\mathbf{H}^{\text{dec}}_{q}\cdot\mathbf{H}^{\text{dec}}_{s})/\tau,\quad\mathcal{L}_{-}=\sum\limits_{s\in S^{-}}{\rm exp}(\mathbf{H}^{\text{dec}}_{q}\cdot\mathbf{H}^{\text{dec}}_{s})/\tau,\\ \mathcal{L}_{\text{ctrs}}&=-\sum\limits_{q\in S^{+}}{\rm log}\left(\frac{\mathcal{L}_{+}}{\mathcal{L}_{+}+\mathcal{L}_{-}}\right),\end{split}$ (3)
where $\tau$ is the temperature parameter, and $\cdot$ denotes inner product.
3.5 Training Loss
The overall loss function is defined as:
$\mathcal{L}=\lambda_{\text{cls}}\mathcal{L}_{\text{cls}}+\lambda_{\text{pos}}\mathcal{L}_{\text{pos}}+\lambda_{\text{ctrs}}\mathcal{L}_{\text{ctrs}},$ (4)
where $\lambda_{\text{cls}}$, $\lambda_{\text{pos}}$, $\lambda_{\text{ctrs}}$ are the coefficients of each loss term, respectively.
4 Experiments
4.1 Datasets and Metrics
RepCountA dataset [13] is currently the largest and most challenging benchmark for the video TRC task  [^1] The RepCountB [13] test subset is proprietary and not publicly available.. It is primarily compiled from fitness videos on YouTube, including a wide range of fitness activities conducted in diverse settings, including homes, gyms, and outdoor environments. This dataset stands out due to its extensive video lengths, significant variations in the average motion cycle, and a higher number of repetitive cycles compared to prior datasets [9, 32, 17, 46]. We use the start and end positions of each action instance $\hat{\mathbf{P}}_{q}$ provided by the annotations, and disregard the class action labels. We train our model on the RepCountA train set and select the best model on the validation set. We report the evaluation results on the test set.
UCFRep dataset [46] is a subset of the UCF101 dataset [35], including fitness videos and daily life videos. Following previous work [13], we do not use the train set but directly test our model on the test set to evaluate the model generalization ability.
Metrics. We compute two commonly used metrics, OBO and MAE  [13, 9, 46], to evaluate the model performance. OBO (Off-By-One count error) measures the percentage that the predicted count is within the GT count ±$\pm 1$ range. MAE (Mean Absolute Error) measures the normalized absolute difference between the predicted and GT counts. Formally,
${\rm OBO}=\frac{1}{M}\sum_{i=1}^{M}\left|N_{i}-\hat{N}_{i}\leq 1\right|,\quad{\rm MAE}=\frac{1}{M}\sum_{i=1}^{M}\frac{\left|N_{i}-\hat{N}_{i}\right|}{N_{i}},$ (5)
where ${N}_{i}$ and $\hat{N}_{i}$ are the predicted and GT counts for the $i^{th}$ test video, respectively, and $M$ is the total number of test videos.
To better evaluate the performance of different models in recognizing actions with varying periods, we expand the evaluation metrics with three variants for the OBO and MAE metrics. We split the test video set into three categories based on the average single action period length: short-, medium-, and long-period test sets. We define videos with an average single action duration of fewer than 30 frames as belonging to the short-period test set, videos with an average action duration longer than 60 frames as the long-period test set, and the remaining videos as belonging to the medium-period test set. We compute the OBO and MAE metrics on each of these sets separately.
4.2 Implementation Details
We implement our approach with different backbone video feature extractors respectively, including TSN [40], I3D [4], and ViT-B [8]. For the Transformer architecture, we employ a 2-layer encoder and a 4-layer decoder, both with 8-head attention mechanisms. The feature dimension is set to $C=512$. We set $Q=40$ queries in our method empirically. For more results with different feature channels $C$ and query numbers $Q$, please refer to the supplementary materials. The length of the video input is set to $T=512$ frames without down-sampling. We utilize the AdamW optimizer [25] with a learning rate of $0.002$, a batch size of $64$, and train the model for $80$ epochs. We set $\lambda_{cls}=1.0$, $\lambda_{pos}=5.0$, $\lambda_{gIoU}=0.4$, $\lambda_{ctrs}=1.0$, confidence threshold $\alpha=0.2$. We provide further implementation details in the supplementary.
4.3 Comparison to State-of-the-arts
Results on RepCountA dataset. We compare our proposed approach to the state-of-the-art methods on the RepCountA [13] dataset. We compare with the SOTA action recognition [10, 24, 23], action segmentation [14], and TRC [9, 46, 13] approaches as explored in TransRAC [13]. We further adapt recent DETR-style action detection approaches [22, 44, 33] for the TRC task. We change their output layers accordingly and train them on RepCountA.
Table 1: Comparison to the state-of-the-arts on RepCountA [13] dataset. We further calculate MAE and OBO metrics for short-, medium-, and long-period test actions. normal-$\dagger$ denotes action detection approaches.
Backbone ${\rm MAE}\downarrow$ ${\rm OBO}\uparrow$ ${\rm MAE_{s}}\downarrow$ ${\rm OBO_{s}}\uparrow$ ${\rm MAE_{m}}\downarrow$ ${\rm OBO_{m}}\uparrow$ ${\rm MAE_{l}}\downarrow$ ${\rm OBO_{l}}\uparrow$
X3D [10] X3D 0.9105 0.1059 - - - - - -
TANet [24] TANet 0.6624 0.0993 - - - - - -
VideoSwinT [23] Swin-T 0.5756 0.1324 - - - - - -
GTRM [14] I3D 0.5267 0.1589 - - - - - -
RepNet [9] ResNet-50 0.5865 0.2450 0.7793 0.0930 0.5893 0.1591 0.4549 0.4062
Zhang et al. [46] 3D-ResNext101 0.8786 0.1554 - - - - - -
TransRAC [13] Swin-T 0.4891 0.2781 0.5789 0.0233 0.4696 0.2955 0.4420 0.4375
TadTR [22] $\dagger$ I3D 1.1314 0.0662 0.8364 0.0233 1.1591 0.0000 1.3106 0.1406
ActionFormer [44] $\dagger$ I3D 0.4990 0.2781 0.4164 0.1628 0.3768 0.3409 0.6385 0.3125
ReAct [33] $\dagger$ TSN 0.4592 0.3509 0.2805 0.1576 0.3037 0.4318 0.6862 0.3906
Ours TSN 0.2809 0.4570 0.2411 0.1628 0.1792 0.5455 0.3776 0.5938
Ours I3D 0.3305 0.4437 0.2421 0.2093 0.2012 0.5227 0.4788 0.5469
Ours ViT-B 0.2622 0.5430 0.2257 0.2558 0.2002 0.5909 0.3294 0.7031
As shown in Tab. 1, our approach significantly outperforms the state-of-the-art methods across actions of varying lengths. Specifically, the similarity-matrix-based methods [9, 13] suffer from quadratic computation complexity. To manage this, they employ a sparse sampling strategy, ensuring reasonable content coverage within a limited temporal context window. However, this approach results in inferior performance for short, rapid action instances, as it tends to overlook the cycles. Conversely, the action detection approaches [22, 44, 33] are primarily developed for detecting action instances specific to particular classes. To adapt them to the TRC task, we modify their output layer to incorporate class-agnostic supervision accordingly. However, relying solely on class-agnostic supervision does not provide them with the ability to dynamically identify repetitive cycles based on the input videos. As a result, their performance on long, slow action instances is generally inferior compared to dedicated TRC approaches. In contrast, our approach effectively balances the detection of actions at various speeds.
Results on UCFRep dataset. We also evaluate the generalization ability of our method. Following previous work [13], we evaluate the model trained on the RepCountA dataset [13] on UCFRep [46] test set. Tab. 2 shows that our approach generally outperforms existing works, with more significant improvements observed in longer-period actions. The results demonstrate the effectiveness and generalization capability of our method.
Table 2: Generalization comparison on UCFRep [46] dataset. We calculate MAE and OBO metrics for short-, medium-, and long-period test actions.
Backbone ${\rm MAE}\downarrow$ ${\rm OBO}\uparrow$ ${\rm MAE_{s}}\downarrow$ ${\rm OBO_{s}}\uparrow$ ${\rm MAE_{m}}\downarrow$ ${\rm OBO_{m}}\uparrow$ ${\rm MAE_{l}}\downarrow$ ${\rm OBO_{l}}\uparrow$
RepNet [9] ResNet-50 0.5336 0.2984 0.6219 0.1739 0.4825 0.3600 0.4996 0.5000
TransRAC [13] Swin-T 0.6180 0.3143 0.6296 0.1951 0.5842 0.4250 0.6784 0.4118
Ours TSN 0.6016 0.2959 0.7069 0.0488 0.5777 0.4250 0.4039 0.5882
Ours I3D 0.5194 0.3980 0.4945 0.2195 0.5865 0.4250 0.4216 0.7647
Ours ViT-B 0.5435 0.4184 0.5657 0.1951 0.4625 0.5500 0.6804 0.6471
Efficiency. Additionally, we evaluate the computational complexity of our method in comparison to state-of-the-art methods. We benchmark all the methods with inference on the RepCountA dataset. For fairness of comparison, we calculate the Floating Point Operations (FLOPs) in the counting module for all the methods, excluding the video feature extraction process. We set the batch size to be $1$ and clip length to be $64$ for each method. Tab. 3 shows that our query-based method significantly reduces the computation load compared to the previous similarity-matric-based methods [9, 13], while also delivering superior accuracy. Importantly, the difference in computational demand is expected to widen with longer input clip lengths. This is because the computational complexity for similarity-matrix-based methods increases quadratically, whereas it grows linearly in our approach. Unfortunately, due to computational limitations, we are unable to obtain the results of RepNet [9] and TransRAC [13] when $T=512$.
Table 3: Comparison of computational complexity.
$T$ Params (M) FLOPs (G)
RepNet [9] 64 20.47 92.40
TransRAC [13] 64 14.43 100.19
Ours 64 18.87 1.44
Ours 512 18.87 6.06
4.4 Qualitative Results
We visualize the predictions of our approach and baseline method TransRAC [13] on the RepCountA [13] dataset to better demonstrate our performance improvement in Fig. 3. TransRAC [13] represents action cycles using density maps, and the final count value is obtained by summing the values in the density map. However, this approach suffers from a lack of interpretability and finally results in miscounting. In contrast, our approach not only produces an accurate final count but also precisely localizes the action start and end positions (colored blocks) in most cases, aligning with the GT labels. It can be seen that our method exhibits robustness to changes in viewpoint, background noise, and sudden interruptions. For instance, in the second case of Fig. 3, we accurately and robustly estimate the time positions even with viewpoint changes as well as a sudden interruption in the middle of the timeline.
In addition, our query-based representation offers excellent interpretability, making it easy to identify the issues in case of failures. For example, Fig. 4 shows a typical failure case. Due to the excessive zooming in, the legs of the human body are truncated, making a large difference in the action motion feature, and resulting in several missed cycle counts.
Figure 3: RepCountA visualization results. Each colored block represents a GT or predicted action instance. TransRAC represents the results by density map, and the final count value is obtained by summing the values in the density map.
Figure 4: Visualization of failure case on RepCountA [13] dataset. Due to the excessive zooming in, the legs of the human body are truncated, making a large difference in the action motion feature, and resulting in several missed cycle counts. Each colored block represents a GT or predicted action instance.
We further illustrate the generalization performance of the proposed method on the unseen UCFRep [46] test set in Fig. 5. We directly apply our trained model and do not use UCFRep training data. Our model still accurately recognizes the action instances and gets the correct count in the challenging cases, indicating robust generalization ability. Specifically, the top case exhibits extreme viewpoint and lighting conditions, while the bottom case contains the action of soccer juggling which is not seen in the training set. We attribute the performance advantage to the proposed DAQ and ICL designs, which empower the model to adaptively adjust the action queries based on the input video features and effectively localize similar (repetitive) action instances, distinguishing them from the background noise actions. Please refer to our supplementary video for more qualitative results which are robust to varying cycle lengths, sudden interruptions, and multiple distracting action classes.
Figure 5: UCFRep visualization results. Each colored block represents a GT or predicted action instance. TransRAC represents the results by density map, and the final count value is obtained by summing the values in the density map. The vertical lines in the labels represent the time points at which the actions begin since only the starting point annotations are provided in UCFRep [46].
4.5 Ablation Study
Effect of DAQ and ICL. We implement two ablated models to study the efficacy of the proposed DAQ and ICL designs. Tab. 4 presents the results on RepCountA [13] using the TSN backbone. In ablation (a), we substitute the proposed DAQ module with a static content query, where the content queries $\mathbf{Q}^{\text{cnt}}$ are also learnable variables, instead of using the content queries filtered out by the encoder and query selection modules. In ablation (b), we eliminate the ICL design among the action queries. The results demonstrate that both DAQ and ICL contribute to performance improvement, particularly in terms of counting medium and long actions.
Table 4: Effect of DAQ and ICL modules on RepCountA [13] dataset. (a) ablates DAQ strategy. (b) ablates the ICL strategy. (c) is our full model.
${\rm MAE}\downarrow$ ${\rm OBO}\uparrow$ ${\rm MAE_{s}}\downarrow$ ${\rm OBO_{s}}\uparrow$ ${\rm MAE_{m}}\downarrow$ ${\rm OBO_{m}}\uparrow$ ${\rm MAE_{l}}\downarrow$ ${\rm OBO_{l}}\uparrow$
(a) w/o DAQ 0.3542 0.4172 0.2624 0.2093 0.2515 0.5227 0.4864 0.4844
(b) w/o ICL 0.4035 0.4040 0.2448 0.2093 0.3106 0.4545 0.5740 0.5000
(c) Ours (full) 0.2809 0.4570 0.2411 0.1628 0.1792 0.5455 0.3776 0.5938
Confidence threshold $\alpha$. We evaluate the impact of different confidence thresholds for the “action of interest” on the final counting results in Fig. 6 (green curve). The results indicate that setting the threshold within the range of $0.2$ to $0.4$ yields similar performance for our method. Our method consistently outperforms TransRAC [13] (blue curve) by a considerable margin.
Figure 6: Results of different confidence thresholds of our method and TransRAC [13]. We depict the MAE (left) and OBO (right) curves of our method (green curve) and the TransRAC (blue curve) approach with regard to different confidence thresholds. The metrics of TransRAC are obtained by binarizing the density map of TransRAC output and then summing to obtain the final count.
5 Conclusion
In conclusion, our study introduces an innovative perspective for the TRC task, which reduces the computational complexity and maintains robustness across varying action periods and video lengths. To tackle the challenge of open-set action categories, we propose DAQ, which improves generalization across different actions. To recognize actions with identical content, we propose ICL, which facilitates learning repetitive action representations and distinguishing them from distractions. Integrating DAQ and ICL, our method can identify contextually similar action instances that are adaptively conditioned on the video content. Experimental results on challenging benchmarks demonstrate the superiority of our approach compared to SOTAs, in terms of both accuracy and efficiency. Furthermore, our approach adeptly balances the demands of diverse action speeds and video durations, establishing a solid foundation for practical implementations in real-world scenarios.

### v2

@@
 Temporal periodicity is a ubiquitous phenomenon in the natural world.
-Temporal Repetition Counting (TRC) aims to accurately measure the number of repetitive action cycles within a given video and holds significant potential for applications ranging from pedestrian detection [29] to fitness monitoring [11].
+Temporal Repetition Counting (TRC) aims to accurately measure the number of repetitive action cycles within a given video and holds significant potential for applications such as fitness monitoring [11] and motion generation [48].
 Pioneer methods  [16, 1, 7, 38, 28, 37, 6] represent time-series video data as one-dimensional signals and employ spectral analysis techniques such as the Fourier transform.
 While suitable for short videos with fixed periodic cycle lengths, these methods struggle to handle real-world scenarios with varying cycle lengths and sudden interruptions.
-Recent studies shift to deep learning-based methods [17, 9, 13, 46] and show promising performances.
+Recent studies shift to deep learning-based methods [17, 9, 13, 46, 18] and show promising performances.
 Notably, most of these methods, such as RepNet [9] and TransRAC [13], utilize a temporal similarity correlation matrix to depict repetitiveness, as illustrated in Fig.
@@
 While these conventional approaches are capable of effectively handling simple motion sequences or those characterized by fixed periodicity, they prove inadequate when confronted with non-stationary motion sequences encountered in real-world scenarios.
-In contrast, deep-learning-based approaches [17, 9, 13, 46] have demonstrated remarkable performance improvements.
+In contrast, deep-learning-based approaches [17, 9, 13, 46, 18] have demonstrated remarkable performance improvements.
 Notably, RepNet [9] and TransRAC [13] leverage temporal similarity matrices of actions to construct models for counting temporal repetitions.
@@
 UCFRep dataset [46] is a subset of the UCF101 dataset [35], including fitness videos and daily life videos.
-Following previous work [13], we do not use the train set but directly test our model on the test set to evaluate the model generalization ability.
+Following previous work [13, 18], we do not use the train set but directly test our model on the test set to evaluate the model generalization ability.
 Metrics.
@@
 Results on RepCountA dataset.
-We compare our proposed approach to the state-of-the-art methods on the RepCountA [13] dataset.
-We compare with the SOTA action recognition [10, 24, 23], action segmentation [14], and TRC [9, 46, 13] approaches as explored in TransRAC [13].
-We further adapt recent DETR-style action detection approaches [22, 44, 33] for the TRC task.
+We compare our proposed approach to the state-of-the-art methods on the RepCountA [13] dataset following previous work [13, 18] in Tab.
+1.
+We compare with the SOTA action recognition [10, 25, 24], action segmentation [14] methods (top block), and TRC [9, 46, 13, 18] approaches (second block).
+We further adapt recent DETR-style action detection approaches [23, 44, 33] for the TRC task (third block).
 We change their output layers accordingly and train them on RepCountA.
 Table 1: Comparison to the state-of-the-arts on RepCountA [13] dataset.
-We further calculate MAE and OBO metrics for short-, medium-, and long-period test actions.
-normal-$\dagger$ denotes action detection approaches.
+We compare with SOTA action recognition/segmentation methods (top block), TRC methods (second block), and action detection methods (third block with normal-$\dagger$).
+We further report MAE and OBO metrics for short-, medium-, and long-period test actions.
 Backbone ${\rm MAE}\downarrow$ ${\rm OBO}\uparrow$ ${\rm MAE_{s}}\downarrow$ ${\rm OBO_{s}}\uparrow$ ${\rm MAE_{m}}\downarrow$ ${\rm OBO_{m}}\uparrow$ ${\rm MAE_{l}}\downarrow$ ${\rm OBO_{l}}\uparrow$
@@
 TransRAC [13] Swin-T 0.4891 0.2781 0.5789 0.0233 0.4696 0.2955 0.4420 0.4375
+Li et al. [18] Swin-T 0.3841 0.3860 - - - - - -
 TadTR [23] $\dagger$ I3D 1.1314 0.0662 0.8364 0.0233 1.1591 0.0000 1.3106 0.1406
@@
 1, our approach significantly outperforms the state-of-the-art methods across actions of varying lengths.
-Specifically, the similarity-matrix-based methods [9, 13] suffer from quadratic computation complexity.
+Specifically, the similarity-matrix-based methods [9, 13, 18] suffer from quadratic computation complexity.
 To manage this, they employ a sparse sampling strategy, ensuring reasonable content coverage within a limited temporal context window.
@@
 We also evaluate the generalization ability of our method.
-Following previous work [13], we evaluate the model trained on the RepCountA dataset [13] on UCFRep [46] test set.
+Following previous work [13, 18], we evaluate the model trained on the RepCountA dataset [13] on UCFRep [46] test set.
 Tab.
@@
 The results demonstrate the effectiveness and generalization capability of our method.
-Table 2: Generalization comparison on UCFRep [46] dataset. We calculate MAE and OBO metrics for short-, medium-, and long-period test actions.
+Table 2: Generalization comparison with SOTA TRC methods on UCFRep [46] dataset. MAE and OBO metrics for short-, medium-, and long-period actions are also reported.
 Backbone ${\rm MAE}\downarrow$ ${\rm OBO}\uparrow$ ${\rm MAE_{s}}\downarrow$ ${\rm OBO_{s}}\uparrow$ ${\rm MAE_{m}}\downarrow$ ${\rm OBO_{m}}\uparrow$ ${\rm MAE_{l}}\downarrow$ ${\rm OBO_{l}}\uparrow$
@@
 TransRAC [13] Swin-T 0.6180 0.3143 0.6296 0.1951 0.5842 0.4250 0.6784 0.4118
+Li et al. [18] 3D-ResNext101 0.5227 0.3500 - - - - - -
 Ours TSN 0.6016 0.2959 0.7069 0.0488 0.5777 0.4250 0.4039 0.5882
@@
 Unfortunately, due to computational limitations, we are unable to obtain the results of RepNet [9] and TransRAC [13] when $T=512$.
-Table 3: Comparison of computational complexity.
+Table 3: Comparison of computational complexity. We benchmark all the methods with inference on the RepCountA dataset.
 $T$ Params (M) FLOPs (G)
@@
 Due to the excessive zooming in, the legs of the human body are truncated, making a large difference in the action motion feature, and resulting in several missed cycle counts.
-Figure 3: RepCountA visualization results.
+Figure 3: Qualitative results on RepCountA dataset.
 Each colored block represents a GT or predicted action instance.
@@
 Please refer to our supplementary video for more qualitative results which are robust to varying cycle lengths, sudden interruptions, and multiple distracting action classes.
-Figure 5: UCFRep visualization results.
+Figure 5: Qualitative results on UCFRep dataset.
 Each colored block represents a GT or predicted action instance.

### v3

@@
-Efficient Action Counting with Dynamic Queries
+
 Abstract
-Temporal repetition counting aims to quantify the repeated action cycles within a video.
-The majority of existing methods rely on the similarity correlation matrix to characterize the repetitiveness of actions, but their scalability is hindered due to the quadratic computational complexity.
+Most existing methods rely on the similarity correlation matrix to characterize the repetitiveness of actions, but their scalability is hindered due to the quadratic computational complexity.
 In this work, we introduce a novel approach that employs an action query representation to localize repeated action cycles with linear computational complexity.
-Based on this representation, we further develop two key components to tackle the essential challenges of temporal repetition counting.
+Based on this representation, we develop two key components to tackle the essential challenges of temporal repetition counting.
 Firstly, to facilitate open-set action counting, we propose the dynamic update scheme on action queries.
@@
 On the challenging RepCountA benchmark, we outperform the state-of-the-art method TransRAC by 26.5% in OBO accuracy, with a 22.7% mean error decrease and 94.1% computational burden reduction.
-Code is available at https://github.com/lizishi/DeTRC.
+Code will be publicly available.
 keywords:
-Temporal repetition counting Video understanding
+Temporal repetition counting, Video understanding
 1 Introduction
@@
 Temporal Repetition Counting (TRC) aims to accurately measure the number of repetitive action cycles within a given video and holds significant potential for applications such as fitness monitoring [11] and motion generation [49].
+Figure 1: Conceptual workflow comparison of (a) similarity-matrix-based methods and (b) our proposed action query-based method.
+(a) Most existing methods use similarity matrices calculated between each frame to detect repetitive actions, resulting in a time complexity of $O(T^{2}C)$, where $T$ denotes video length and $C$ denotes feature dimension.
+(b) On the other hand, our method employs action queries to represent each action cycle and estimates their classes and temporal locations.
+Each query is classified into either “repetitive actions” or “others ($\emptyset$)” class labels.
+The temporal location is expressed by the midpoint and duration on the timeline.
+During training, we perform bipartite matching to uniquely associate a prediction to a GT action cycle, taking into account both class labels and temporal locations.
+Predictions with no match should yield a $\emptyset$ class prediction, indicating that these are not repetitive actions.
+This novel formulation reduces the complexity from quadratic to linear as $O(TC)$.
 Pioneer methods  [17, 1, 7, 39, 29, 38, 6] represent time-series video data as one-dimensional signals and employ spectral analysis techniques such as the Fourier transform.
@@
 Recent progress in action detection [34, 24] introduces an efficient representation of action periods by associating each action instance with an action query, similar to DETR [3].
-Inspired by this, we propose to formulate the TRC problem as a set prediction task where the goal is to detect every action cycle by representing it as an action query.
-This novel formulation reduces the complexity from quadratic to linear and enables counting long videos with varying action periods, as illustrated in Fig.
+Inspired by this, we propose to formulate the TRC problem as a set prediction task where the goal is to detect every action cycle by representing it as an action query, as illustrated in Fig.
 1 (b).
+Based on this query representation, we use a Transformer encoder-decoder network [50] to detect repetitive action instances and their temporal positions, defined by their midpoints and durations.
+This novel formulation reduces the complexity from quadratic to linear  [^1] In implementation, we employ deformable attention modules as proposed in DeformableDETR [50].
+and enables counting long videos with varying action periods.
 However, directly applying the action detection approaches [24, 45, 34] to the TRC problem proves inadequate (Tab.
@@
     TRC requires recognizing action instances with identical content, while detection does not.
-Figure 1: Comparison of (a) similarity-matrix-based methods and (b) our proposed action query-based method.
-Unlike the similarity-matrix-based methods, which exhibit a quadratic growth $O(T^{2}C)$ in computational complexity as the length of the input video $T$ increases, our approach is characterized by a linear growth $O(TC)$ in complexity, where $C$ denotes the feature dimension of the video frames.
-This distinct advantage allows us to effectively process long videos that contain varying durations of action sequences.
 As a result, approaching TRC as a simple action detection task results in inferior performance as shown in Sec.
@@
 In contrast, we propose two novel strategies to address these challenges and redefine the framework for TRC.
-In response to the first challenge, we propose Dynamic Action Query (DAQ), which adaptively updates the action query using distilled content features from the video encoder.
-This mechanism allows the decoder to attend to the “action of interest” based on the input video contents in a dynamic and contextually aware manner, without the need for manual definition.
-To tackle the second challenge, we propose Inter-query Contrastive Learning (ICL).
-It enforces that the repetitive action cycles are grouped together in the learned representation space while being separated apart from non-repetitive (e.g.
-background or distracting) video content.
+In response to the first challenge, we propose Dynamic Action Query (DAQ), which adaptively updates the action query using content features extracted from the video (Fig.
+2).
+This mechanism allows the decoder to attend to the repetitive actions based on the input video contents in a dynamic and contextually aware manner, without the need for manual definition.
+To tackle the second challenge, we define two action classes: “repetitive actions”, which are dynamically defined by DAQ, and the “others ($\emptyset$)” class.
+We further propose Inter-query Contrastive Learning (ICL) which ensures that primary repetitive action cycles are grouped together in the learned representation space while being separated from other distractors, such as background noise.
 The integration of two core components (DAQ and ICL) ensures that the action instances are identified adaptively based on the video content and their contextual similarity.
@@
   2. 2.
-    We propose Dynamic Action Query to guide the model to focus on the action of interest and improve generalization ability across different actions.
+    We propose Dynamic Action Query to guide the model to focus on the repetitive actions contextually defined by the video content thereby improving generalization ability across different actions.
   3. 3.
-    We introduce Inter-query Contrastive Learning to facilitate learning repetitive action representations and to distinguish these from distractions.
+    We introduce Inter-query Contrastive Learning to facilitate learning primary repetitive action representations and to distinguish them from distractions.
   4. 4.
@@
 3 Method
+3.1 Preliminary
+DETR [3] is a pioneering object detection framework that builds upon the Transformer encoder-decoder architecture [40].
+The overall DETR architecture [3] consists of three main components: a backbone to extract image features, an encoder-decoder Transformer, and the detection heads, i.e.
+feed-forward network (FFN) that makes the final detection prediction.
+The main features of DETR are the conjunction of a bipartite matching loss and transformers with (non-autoregressive) parallel decoding.
+The bipartite matching loss is a set-based Hungarian loss during training, which uniquely assigns a prediction to a GT object, and is invariant to a permutation of predicted objects.
+This design enables DETR to perform parallel processing and predict all objects simultaneously.
+We briefly review the workflow as follows.
+Given the input image feature maps extracted by a CNN backbone, e.g.
+ResNet [13], DETR exploits a standard Transformer encoder-decoder network to transform the feature maps to be features of a set of object queries.
+An FFN and a linear projection are added on top of the object query features as the detection heads.
+The FFN acts as the regression branch that predicts the bounding box coordinates, i.e.
+box center coordinates, box height and width.
+The linear projection acts as the classification branch to produce the classification results, i.e.
+$\mathtt{object}$ vs.
+$\mathtt{no\,object}\,(\emptyset)$.
+The $\emptyset$ class is used to represent that no object is detected, playing a similar role to the “background” class in the standard object detection approaches.
+DETR infers a fixed-size set of $N$ predictions in a single pass, where $N$ is set to be significantly larger than the typical number of objects in an image.
+During training, a Hungarian loss produces an optimal bipartite matching between predicted and GT objects and then optimizes the object position-specific losses.
+The matching procedure takes into account both the class prediction and the similarity of predicted and GT boxes and finds one-to-one matching for direct set prediction without duplicates.
+Figure 2: Overview and detailed architecture design of our method.
+(Left) We utilize a DETR-inspired framework to predict the number of repetitive action cycles in an RGB video sequence.
+The architecture consists of a backbone network $\Phi(\cdot)$ for feature extraction, an encoder-decoder Transformer for processing query features, and prediction heads for action classification.
+By counting the queries classified as “repetitive actions”, we get the total count $N$.
+(Right) The detailed workflow within the black box on the left.
+Given the video features $\mathbf{F}$ as input, the encoder $\mathcal{E}(\cdot)$ produces two sets of query features, i.e.
+action queries and position queries.
+A query selection module then screens these queries, retaining only the most relevant ones for forwarding to the decoder $\mathcal{D}(\cdot)$.
+This process utilizes action classification results from prediction heads positioned after the encoder.
+The decoder adopts the Dynamic Action Query (DAQ) strategy, utilizing the selected action queries as priors, while incorporating a learnable position query and attending to the encoder’s output.
+The output embeddings from the decoder are then passed to their corresponding prediction head, which estimates the class label $\mathbf{A}$ and temporal location $\mathbf{P}$.
+During training, we employ bipartite matching to uniquely pair each prediction with a GT action cycle $\hat{\mathbf{P}}$, considering both class labels and temporal locations.
+Predictions that fail to match should yield $\emptyset$ class, signifying that they do not represent repetitive actions.
+The DAQ strategy addresses the open-set problem by contextually defining “repetitive actions” based on video features.
+To further distinguish these actions from $\emptyset$ distractions, we propose Inter-query Contrastive Learning (ICL).
+By integrating DAQ and ICL, our method effectively identifies contextually similar actions and excludes distractions, making it well-suited for tackling the challenges of the TRC problem.
 3.2 Overview
-Given an RGB video sequence $\mathbf{V}$ with $T$ frames, The TRC task aims to predict an integer $N$ indicating the number of estimated repetitive action cycles.
-Our model consists of a video backbone network $\Phi(\cdot)$, a Transformer network with encoder $\mathcal{E}(\cdot)$, decoder $\mathcal{D}(\cdot)$, and multiple prediction heads as shown in Fig.
+Given an RGB video sequence with $T$ frames, the TRC task aims to predict an integer $N$ indicating the number of detected repetitive action cycles.
+Drawing inspiration from DETR [3], we streamline the problem as a set detection task in the temporal domain.
+The overall framework consists of three main components: a backbone network $\Phi(\cdot)$, an encoder-decoder Transformer, and prediction heads, as shown in Fig.
 2.
+The backbone extracts video features $\mathbf{F}$, which is transformed to features of a set of action queries by an encoder $\mathcal{E}(\cdot)$ and a decoder $\mathcal{D}(\cdot)$.
+Two prediction heads are added on top of the query features.
+The prediction heads consist of an action head and a position head.
+The action head classifies the query, while the position head estimates its temporal location.
+Building on previous enhancements in DETR models [23], we introduce two distinct types of queries to enhance the query-to-feature similarity and accelerate convergence: action queries and position queries.
+Action queries are designed to capture features pertinent to action classification, whereas position queries concentrate on the temporal dimensions of an action.
+As depicted in the detailed diagram on the right side of Fig.
+2, two specialized prediction heads (i.e.
+action head and position head) are tasked with processing these respective sets of queries to generate predictions.
+To further accelerate convergence, we introduce a post-encoder query selection module that acts as an initial sieve, filtering out the majority of queries unlikely to represent repetitive actions before forwarding the most promising ones to the decoder.
+This enhances the efficiency and focus of the subsequent prediction process.
+Considering the two challenges of the TRC tasks discussed in Sec.
+1, we propose Dynamic Action Query (DAQ) strategy to address the open-set problem.
+DAQ directly uses dynamically updated action query features from the encoder to initialize the action query of the decoder.
+This strategy defines what “repetitive actions” should be detected based on the video content features in a dynamic and contextually aware manner.
+To tackle the second challenge, we define two action classes: “repetitive actions” and the “others ($\emptyset$)” class.
+We further propose Inter-query Contrastive Learning (ICL) to cluster repetitive action queries into a positive action set ($S^{+}$) and group the other queries to a negative set ($S^{-}$) in the feature space.
+Integrating DAQ and ICL allows our method to identify contextually similar action instances that are adaptively based on the video content, and exclude other distracting actions at the same time.
+Similar to DETR [3], during training, we employ bipartite matching which uniquely assigns a prediction to a GT action cycle.
+The matching procedure takes into account both the class prediction and the similarity of predicted and GT temporal locations.
+We then optimize the action cycle-specific losses.
+In inference, our method produces a fixed-size set of $Q$ predictions in a single pass, where $Q$ is set to be significantly larger than the typical number of action cycles in a $T$-frame video sequence.
+By counting the queries classified as “repetitive actions”, we get the final total count value $N$.
+In the following, we will introduce the network architecture design in Sec. 3.3, the DAQ and ICL modules in Sec. 3.4 and Sec. 3.5, and the model training in Sec. 3.6.
+3.3 Model Architecture
+Backbone.
 The backbone network $\Phi(\cdot)$ takes a sequence of $T$ video frames as input and extracts feature vectors $\mathbf{F}\in\mathbb{R}^{T\times C}$ for each frame, where $C$ denotes the feature dimension.
-In practice, we utilize the commonly-used backbone works, e.g.
-TSN [40], and freeze the backbone parameters.
-Figure 2: Framework overview.
-Given a video input, we first extract video features $\mathbf{F}$ with the backbone network Φ()normal-$\Phi(\cdot)$.
-Then we leverage a Transformer encoder-decoder network to process the features $\mathbf{F}$ to detect each action cycle instance by predicting its action class label $\mathbf{C}^{\text{cls}}$ and its time position $\mathbf{P}$.
-To tackle the challenge of open-set action categories, we propose Dynamic Action Query (DAQ), allowing the method to attend to the “action of interest” based on the input video contents in a dynamic and contextually aware manner.
-To recognize actions with identical content, we propose Inter-query Contrastive Learning (ICL) to group the “action of interest” and be separated apart from noise actions (e.g.
-background or distractions).
-Integrating DAQ and ICL, our method can identify contextually similar action instances that are adaptively conditioned on the video content, without manual action class definition.
-By counting the queries classified as the “action of interest” with classification confidence score $\mathbf{C}^{\text{conf}}$ exceeding the threshold $\alpha$, we finally get the total count value $N$.
-We then design a Transformer encoder-decoder architecture to process the features $\mathbf{F}$ to detect each action cycle instance and obtain the final action cycle count value $N$.
-This requires the model to discern each detected action cycle class $\mathbf{C}^{\text{cls}}$ and its time position $\mathbf{P}$.
-To achieve this, we introduce the concept of action queries within the decoder $\mathcal{D}(\cdot)$, which are comprised of content queries $\mathbf{Q}^{\text{cnt}}\in\mathbb{R}^{Q\times C}$ and position queries $\mathbf{Q}^{\text{pos}}\in\mathbb{R}^{Q\times C}$, where $Q$ represents the number of queries.
-Refer to Fig.
-2 for illustration.
-The content queries $\mathbf{Q}^{\text{cnt}}$ are responsible for capturing the action category, while the position queries $\mathbf{Q}^{\text{pos}}$ handle the action interval estimation.
-We use two prediction heads including a classification head and a position head after the decoder to predict the class label $\mathbf{C}^{\text{cls}}$ and time position $\mathbf{P}$ for each query, respectively.
-To tackle the challenge of open-set action categories, we define $K=2$ classes, namely “action of interest” and “null ($\emptyset$)”.
-We then propose the Dynamic Action Query (DAQ) strategy which adaptively updates the content query $\mathbf{Q}^{\text{cnt}}$ of the decoder using distilled content features that are selected from the encoder $\mathcal{E}(\cdot)$.
-This mechanism eliminates the need for a manual definition of the specific semantics of “action of interest” and allows the decoder to attend to the “action of interest” conditioned on the input video contents in a dynamic and contextually-aware manner.
-Nevertheless, videos may contain distracting factors such as background elements or other noise actions, necessitating that the detected “action of interest” instances exhibit identical or similar contents.
-To this end, we propose Inter-query Contrastive Learning (ICL) to cluster content queries into positive action set ($S^{+}$) and negative null set ($S^{-}$) based on their predicted class label $\mathbf{C}^{\text{cls}}$.
-Integrating DAQ and ICL allows our method to identify contextually similar action instances that are adaptively based on the video content, and exclude other distracting actions at the same time.
-Finally, by counting the queries classified as the “action of interest” with classification confidence score $\mathbf{C}^{\text{conf}}$ exceeding a threshold $\alpha$, we get the total count value $N$.
-3.2 Model Architecture
 Encoder.
-The encoder $\mathcal{E}(\cdot)$ is a classical Transformer [39] architecture which processes the video features $\mathbf{F}$ extracted by the backbone network.
-The output features $\mathbf{H}^{\text{enc}}$ of $\mathcal{E}(\cdot)$ will then be fed into a query selection module to select $Q$ queries to be the initial content queries $\mathbf{Q}^{\text{cnt}}$ for the decoder.
-For clarity, we will detail the query selection module later in this section.
-At now, it can be considered as a rough filter that selects top-$Q$ relevant features as the input queries for decoder $\mathcal{D}(\cdot)$.
+The encoder $\mathcal{E}(\cdot)$ is a classical Transformer [40] architecture which has ${L}^{\text{enc}}$ standard encoder layers.
+The encoder transforms the video features $\mathbf{F}$ into two distinct query types: action queries $\tilde{\mathbf{E}}^{\text{act}}\in\mathbb{R}^{T\times C}$ and position queries $\tilde{\mathbf{E}}^{\text{pos}}\in\mathbb{R}^{T\times C}$, as shown in Fig.
+2 (right).
+This design enhances the query-to-feature similarity and accelerates convergence, where action queries capture features pertinent to action classification.
+On the other hand, position queries concentrate on the temporal dimensions of an action.
+We defer to the supplementary material the detailed encoder architecture.
+Query selection.
+We add a query selection module before passing the encoder output query features to the decoder.
+We route the query features outputted by the encoder to two respective prediction heads (described later in this section), which independently decode them into predictions for action class and temporal location, as depicted on the right side of Fig.
+2.
+The action head processes the action queries $\tilde{\mathbf{E}}^{\text{act}}$ to estimate the action class $\tilde{\mathbf{A}}$ for each query, while the position head decodes the position queries $\tilde{\mathbf{E}}^{\text{pos}}$ to their temporal locations $\tilde{\mathbf{P}}$.
+Based on the prediction results, we discard query features classified as “others ($\emptyset$)”, retaining only those identified as “repetitive actions”.
+For the remaining queries, we rank them based on their classification confidence and preserve only the top $Q$ high-confidence queries.
+The others are discarded.
+After this selection process, both the action query features $\mathbf{E}^{\text{act}}$ and position query features $\mathbf{E}^{\text{pos}}$ have a dimension of $Q\times C$.
+Additional details will be provided when we discuss the prediction head later in this section.
 Decoder.
-The decoder $\mathcal{D}(\cdot)$ is also a classical Transformer [39] architecture, which takes the encoder features $\mathbf{H}^{\text{enc}}$ and two types of action queries as input and outputs a processed feature $\mathbf{H}^{\text{dec}}\in\mathbb{R}^{Q\times C}$.
-The action queries are constructed by content queries $\mathbf{Q}^{\text{cnt}}\in\mathbb{R}^{Q\times C}$ and position queries $\mathbf{Q}^{\text{pos}}\in\mathbb{R}^{Q\times C}$, where $\mathbf{Q}^{\text{cnt}}$ and $\mathbf{Q}^{\text{pos}}$ are designed for representing the content features and the time position of each action cycle, respectively.
-The content queries $\mathbf{Q}^{\text{cnt}}$ are directly initialized from the distilled content features selected from the encoder $\mathcal{E}(\cdot)$, i.e.
-the proposed DAQ strategy, which will be detailed in Sec.
-3.3.
-The position queries $\mathbf{Q}^{\text{pos}}$ are initialized as learnable parameters and transformed to the initial time position $\mathbf{P}^{\rm init}=(\mathbf{m}^{\rm init},\mathbf{d}^{\rm init})\in\mathbb{R}^{Q\times 2}$ with a linear layer before feeding into the decoder $\mathcal{D}(\cdot)$, where we propose to use the midpoint $\mathbf{m}$ and the period duration $\mathbf{d}$ to denote each action time position.
-For clarity, we omit the linear layer in Fig.
-2.
-Prediction heads. The prediction heads consist of a classification head network and a position head network, which are both MLPs.
-The classification head takes the output features of the decoder $\mathbf{H}^{\text{dec}}$ as input and predicts the action class label for each content query.
-Given that the action category is an open set and not preset, we set the number of classes $K$ to be 2 in the classification head, indicating whether it is an “action of interest” or a “null ($\emptyset$)” class.
-The classification head employs the Softmax function to calculate the probabilities $\mathbf{C}^{\text{cls}}$ of $K$ categories for each action query.
-We then designate the category with the highest probability as the predicted class label for the respective action query.
-Subsequently, we regard this probability value as the estimated confidence score $\mathbf{C}^{\text{conf}}$.
-We define an action query instance as “action of interest” when its confidence score $\mathbf{C}^{\text{conf}}$ is higher than the threshold $\alpha$.
-We apply binary Cross-entropy loss to supervise the predicted class labels $\mathbf{C}^{\text{cls}}$ for each action instance, aiming to guide the model in effectively discriminating the action instances of interest from others:
-$\mathcal{L}_{\text{cls}}=\sum\limits_{q=1}^{Q}{\rm CrossEntropyLoss}\left(\mathbf{C}^{\text{cls}}_{q},\tilde{\mathbf{C}}^{\text{cls}}_{q}\right)=\sum\limits_{q=1}^{Q}\sum\limits_{i=1}^{K}-\tilde{c}_{q,i}{\rm log}\left(c_{q,i}\right),$ (1)
-where $\mathbf{C}^{\text{cls}}_{q}=[c_{q,1},c_{q,2},...,c_{q,K}]\in\mathbb{R}^{K}$ represents the predicted class label for $q^{th}$ action query. $\tilde{}$ denotes the GT label.
-The position head estimates the residual of the midpoint and duration based on the decoder features $\mathbf{H}^{\text{dec}}$ for each query, and adds to the initial time position $\mathbf{P}^{\rm init}$ to get the final estimated time position.
-We supervise the predicted time positions with $L_{1}$ loss and gIoU [31] loss:
-$\begin{split}\mathcal{L}_{\text{pos}}&=\sum\limits_{q=1}^{Q}\left\|\mathbf{P}_{q}-\tilde{\mathbf{P}}_{q}\right\|+\lambda_{\text{gIoU}}\sum\limits_{q=1}^{Q}\left(1-{\rm gIoU}(\mathbf{P}_{q},\tilde{\mathbf{P}}_{q})\right),\end{split}$ (2)
-where $\mathbf{P}_{q}=(\mathbf{m}_{q},\mathbf{d}_{q})$ represents the predicted action instance position for the $q^{th}$ query. $\tilde{}$ stands for GT label.
-Note that the $Q$ predicted instances from the prediction heads are in a random order.
-Therefore, before applying the two loss terms, we have to establish a one-to-one correspondence between the disordered predicted instances and the ground-truth (GT) labels.
-Concretely, we employ the Hungarian matching algorithm [15] and the matching process considers two criteria: 1) the action class label is the “action of interest”, and 2) the time position of the action must align with a corresponding ground-truth time position label.
-A matching cost matrix is constructed based on these criteria.
-For the first criterion, the Negative Log-Likelihood (NLL) loss is employed to assess classification accuracy.
-For the second aspect, both Intersection over Union (IoU) and $L_{1}$ distance are utilized to compute costs, measuring the alignment between the estimated time positions and GT labels.
-The matching cost matrix is of dimensions $Q\times N_{gt}$, where $N_{gt}$ represents the number of GT action instances.
-Utilizing the Hungarian matching algorithm [15] based on this cost matrix, we derive the prediction-to-GT matching results.
-Query selection.
-As we mentioned above, videos may contain multiple action categories or other distracting factors such as background elements, therefore, we design a query selection module after the encoder $\mathcal{E}(\cdot)$ to filter the “action of interest” coarsely.
-To achieve this, we reuse the two prediction heads to predict the class labels and time positions for the output features $\mathbf{H}^{\text{enc}}$ of the encoder $\mathcal{E}(\cdot)$.
-Based on their predicted confidence scores $\mathbf{C}^{\text{conf}}$, the query selection module selects top-$Q$ queries with high confidence that are classified as the “action of interest” to be the initial content queries $\mathbf{Q}^{\text{cnt}}$ for the decoder.
-For clarity, we omit the encoder prediction heads in Fig.
-2.
-Note that the two prediction head networks are shared between the encoder and decoder.
-In addition, we apply corresponding supervision using $\mathcal{L}_{\text{cls}}$ and $\mathcal{L}_{\text{pos}}$ to the encoder prediction heads.
-Additional experimental results of the encoder prediction heads can be found in the supplementary materials.
+The decoder $\mathcal{D}(\cdot)$ is also a classical Transformer [40] architecture, consisting of ${L}^{\text{dec}}$ standard decoder layers.
+It processes a set of action queries $\mathbf{Q}^{\text{act}}\in\mathbb{R}^{Q\times C}$ and position queries $\mathbf{Q}^{\text{pos}}\in\mathbb{R}^{Q\times C}$, while simultaneously attending to the direct output (i.e.
+$\tilde{\mathbf{E}}^{\text{act}}$ and $\tilde{\mathbf{E}}^{\text{pos}}$) from the encoder, as depicted in Fig.
+2 (right).
+Following the practice in DETR [3], these inputs are decoded in parallel across each decoder layer and transformed into the corresponding query features $\mathbf{D}^{\text{act}}\in\mathbb{R}^{Q\times C}$ and $\mathbf{D}^{\text{pos}}\in\mathbb{R}^{Q\times C}$.
+The input position queries $\mathbf{Q}^{\text{pos}}$ are initialized as learnable parameters.
+To address the open-set challenge in the TRC task, we introduce the DAQ strategy, which initializes the decoder input action queries using the selected encoder embeddings $\mathbf{E}^{\text{act}}$, i.e.
+$\mathbf{Q}^{\text{act}}=\mathbf{E}^{\text{act}}$.
+The details of the DAQ strategy will be elaborated in Sec.
+3.4.
+Using self- and encoder-decoder attention over these embeddings, our method globally reasons selected video frame features together while being able to use the whole frame-wise video feature as context.
+The two sets of $Q$ queries are then independently transformed into corresponding prediction results, i.e.
+an action head decodes $\mathbf{D}^{\text{act}}$ into the action class prediction $\mathbf{A}$, and a position head decodes $\mathbf{D}^{\text{pos}}$ into the temporal location prediction $\mathbf{P}$.
+We detail the two heads in the following subsection.
+Prediction heads.
+The prediction heads are designed to make the final predictions using the features processed by the encoder-decoder.
+They comprise two networks: an action head and a position head, both of which are Multi-Layer Perceptrons (MLPs).
+The action head processes the output action queries from the decoder $\mathbf{D}^{\text{act}}$ to estimate the action class $\mathbf{A}\in\mathbb{R}^{Q\times 1}$ for each query using a softmax function.
+Each element of $\mathbf{A}$ represents the probability that the corresponding query is a “repetitive action”, with values spanning from 0 to 1.
+In practice, a threshold value, $\alpha$, is used to categorize the queries: values greater than $\alpha$ classify a query as a “repetitive action” while values less than or equal to $\alpha$ categorize it as “others ($\emptyset$)”.
+Recall that in the query selection module, we rank the queries based on their estimated probabilities $\tilde{\mathbf{A}}\in\mathbb{R}^{T\times 1}$, which serve as confidence scores.
+We then select the top $Q$ queries with the highest confidence scores according to the ranking.
+The position head estimates the temporal location $\mathbf{P}=(\mathbf{m},\mathbf{d})\in\mathbb{R}^{Q\times 2}$ for each query based on the output position queries from the decoder $\mathbf{D}^{\text{pos}}$.
+We use the midpoint $\mathbf{m}\in\mathbb{R}^{Q}$ and the duration $\mathbf{d}\in\mathbb{R}^{Q}$ to denote the temporal location of an action cycle.
+Bipartite matching.
+Following the practice of DETR [3], our method produces a fixed-size set of $Q$ predictions in a single pass, where $Q$ is set to be much larger than the typical number of action cycles in a $T$-frame video sequence.
+Note that the $Q$ predictions are in a random order.
+One of the main difficulties of training is to match the predicted action cycles (classes and temporal locations) to the GT action cycles, therefore, optimizing action cycle-specific losses, e.g.
+classification loss, temporal position loss, etc.
+Let us denote GT repetitive action cycle labels by $\hat{\mathbf{Y}}=(\hat{\mathbf{A}},\hat{\mathbf{P}})$.
+$\hat{\mathbf{A}}=\mathbf{1}\in\mathbb{R}^{\hat{N}\times 1}$ denotes the $\hat{N}$ repetitive action class labels, where we assign the value 1 to the “repetitive actions”.
+$\hat{\mathbf{P}}\in\mathbb{R}^{\hat{N}\times 2}$ denotes the GT temporal positions of the $\hat{N}$ repetitive action, represented by $\hat{N}$ sets of midpoints and duration.
+Similarly, we denote the predicted repetitive actions by $\mathbf{Y}=(\mathbf{A},\mathbf{P})$, a set of size $Q$ predictions.
+We set $Q$ much larger than $\hat{N}$ empirically.
+We consider $\hat{\mathbf{Y}}$ also a set of size $Q$ padded with “others ($\emptyset$)”, i.e.
+we assign the value 0 to denote $\emptyset$ class.
+To find a bipartite matching between these two sets $\hat{\mathbf{Y}}$ and $\mathbf{Y}$, we search for a permutation of $Q$ elements $\sigma\in\mathfrak{S}_{Q}$ with the lowest cost:
+$\hat{\sigma}=\arg\min_{\sigma\in\mathfrak{S}_{Q}}\sum_{i=1}^{Q}\mathcal{L}_{\text{match}}(\hat{\mathbf{Y}}_{i},\mathbf{Y}_{\sigma(i)}),$ (1)
+where $\mathcal{L}_{\text{match}}$ is a pair-wise matching cost between GT action cycle $\hat{\mathbf{Y}}_{i}$ and a prediction cycle with index $\sigma(i)$.
+The matching cost takes into account both the action classification result and the similarity of the predicted temporal locations and the GT temporal location.
+Each element $i$ of the GT action cycle set can be seen as $\hat{\mathbf{Y}}_{i}=(\hat{\mathbf{A}}_{i},\hat{\mathbf{P}}_{i})$ where $\hat{\mathbf{A}}_{i}$ is the target class label (1 for “repetitive actions”, and 0 for $\emptyset$) and $\hat{\mathbf{P}}_{i}\in\mathbb{R}^{2}$ is a vector that denotes the midpoint time position and lasting duration of a GT action cycle.
+For the prediction with index $\sigma(i)$, we define probability of class $\hat{\mathbf{A}}_{i}$ as $p_{\sigma(i)}(\hat{\mathbf{A}}_{i})$.
+Then we can define the matching cost as
+$\begin{split}\mathcal{L}_{\text{match}}(\hat{\mathbf{Y}}_{i},\mathbf{Y}_{\sigma(i)})&=-\mathbbm{1}_{\{\hat{\mathbf{A}}_{i}\neq\emptyset\}}p_{\sigma(i)}(\hat{\mathbf{A}}_{i})\\ &+\mathbbm{1}_{\{\hat{\mathbf{A}}_{i}\neq\emptyset\}}\mathcal{L}_{\text{pos}}(\hat{\mathbf{P}}_{i},\mathbf{P}_{\sigma(i)}),\end{split}$ (2)
+where $\mathbbm{1}$ is an indicator function.
+To measure the similarity of the predicted temporal locations and the GT temporal location, we define $\mathcal{L}_{\text{pos}}$ using the linear combination of the $L_{1}$ distance and the generalized Intersection over Union (IoU) loss [32].
+Overall, the position loss is defined as
+$\begin{split}\mathcal{L}_{\text{pos}}(\hat{\mathbf{P}}_{i},\mathbf{P}_{\sigma(i)})&=\lambda_{\text{L1}}\|\hat{\mathbf{P}}_{i}-\mathbf{P}_{\sigma(i)}\|_{1}\\ &+\lambda_{\text{gIoU}}\mathcal{L}_{\text{gIoU}}(\hat{\mathbf{P}}_{i},\mathbf{P}_{\sigma(i)}),\end{split}$ (3)
+where $\lambda_{\text{gIoU}},\lambda_{\text{L1}}\in\mathbb{R}$ are hyperparamters.
+By employing the Hungarian matching algorithm [16] to optimize Eq.
+1, we can achieve the final optimal matching $\hat{\sigma}$ which uniquely assigns a prediction to a GT action cycle, finding one-to-one matching without duplicates.
+Notice that the matching cost between a repetitive action instance and $\emptyset$ doesn’t depend on the prediction, which means that in that case the cost is a constant.
+Following this step, we can apply the corresponding losses related to classification and temporal location prediction.
+We define a Hungarian loss for all pairs matched in the previous step to supervise both classification and temporal location predictions, i.e.
+a linear combination of a negative log-likelihood loss for class prediction and a temporal position loss defined in Eq.
+3:
+$\begin{split}\mathcal{L}_{\text{Hungarian}}(\hat{\mathbf{Y}},\mathbf{Y})&=\sum\limits_{i=1}^{Q}\left[-\text{log}\,{p_{\hat{\sigma}(i)}}{(\hat{\mathbf{A}}_{i})}\right.\\ &+\left.\mathbbm{1}_{\{\hat{\mathbf{A}}_{i}\neq\emptyset\}}\mathcal{L}_{\text{pos}}(\hat{\mathbf{P}}_{i},\mathbf{P}_{\hat{\sigma}(i)})\right],\end{split}$ (4)
+where $\hat{\sigma}$ is the optimal assignment computed in the previous bipartite matching step, i.e. Eq. 1.
 3.4 Dynamic Action Query
@@
 1, the TRC problem requires recognizing open-set action instances depending on the video content, where the action category is not predefined.
-Therefore, we propose the Dynamic Action Query strategy, which adaptively updates the content query $\mathbf{Q}^{\text{cnt}}$ using distilled content features from the encoder.
-Simple yet effective, the DAQ strategy directly initializes the input content queries $\mathbf{Q}^{\text{cnt}}$ of the decoder using the $Q$ content queries filtered by the query selection module, as shown in Fig.
+Therefore, we propose the Dynamic Action Query strategy, which adaptively updates the action query $\mathbf{Q}^{\text{act}}$ using the selected encoder action query features $\mathbf{E}^{\text{act}}$, i.e.
+$\mathbf{Q}^{\text{act}}=\mathbf{E}^{\text{act}}$, as shown in Fig.
 2.
-This mechanism provides priors for the decoder that allows it to attend to the “action of interest” conditioned on the input video contents in a dynamic and contextually aware manner.
-In this way, we avoid the need for the manual definition of the action categories, thereby improving the models’ generalization capability.
-Furthermore, we extensively explore different methods for initializing the queries in the decoder $\mathcal{D}(\cdot)$ in the supplementary material, confirming the DAQ strategy as the most effective.
+The DAQ strategy, simple yet effective, provides priors for the decoder to dynamically and contextually attend to “repetitive actions” based on the input video content, eliminating the need for manually defining action categories and thereby enhancing the model’s generalization capability.
+This mechanism ensures that the cross-attention module in the decoder pools features from the entire video while focusing on relevant local repetitive features.
+Such dynamic adjustment allows the model to concentrate on pertinent “repetitive actions”, offering a deeper understanding of action queries and improving training convergence.
+In summary, the DAQ approach not only circumvents the need for manual action category definitions but also leverages dynamic, contextually aware priors to enhance both the focus and generalization capabilities of the model.
+Our supplementary material further explores various methods for initializing the queries in the decoder $\mathcal{D}(\cdot)$, confirming that the DAQ strategy is the most effective.
 3.5 Inter-query Contrastive Learning
-Since the input video may contain multiple action categories and other distractors such as the background motion, another unique challenge to the TRC task is to recognize action instances with identical content.
+Since the input video may contain other distractors such as the background motion  [^2] We assume that each video contains only a primary repetitive action type, which is also the case for the public dataset., another unique challenge to the TRC task is to recognize action instances with identical content.
 This requires that the actions of interest we classify exhibit similarity in their motion patterns, while other action queries should have dissimilar representations.
 To tackle this challenge, we propose Inter-query Contrastive Learning to distinguish the action queries.
-Intuitively, we hope to partition the content queries into $K=2$ main categories through contrastive learning.
-One category comprises the “actions of interest”, with similar representations, while the other consists of null instances.
-The latter may encompass interfering actions or background elements, with distinct representations from those of the actions of interest.
-Concretely, we designate the $Q$ decoder features $\mathbf{H}^{\text{dec}}$ features classified into two sets according to the predicted class labels $\mathbf{C}^{\text{cls}}$, as shown in Fig.
-2.
-The features that are classified as actions of interest form the positive set $S^{+}$, while the remaining features form the negative set $S^{-}$.
-Then we apply contrastive learning using InfoNCE loss [12] over the feature space:
-$\begin{split}\mathcal{L}_{+}&=\sum\limits_{s\in S^{+},s\neq q}{\rm exp}(\mathbf{H}^{\text{dec}}_{q}\cdot\mathbf{H}^{\text{dec}}_{s})/\tau,\quad\mathcal{L}_{-}=\sum\limits_{s\in S^{-}}{\rm exp}(\mathbf{H}^{\text{dec}}_{q}\cdot\mathbf{H}^{\text{dec}}_{s})/\tau,\\ \mathcal{L}_{\text{ctrs}}&=-\sum\limits_{q\in S^{+}}{\rm log}\left(\frac{\mathcal{L}_{+}}{\mathcal{L}_{+}+\mathcal{L}_{-}}\right),\end{split}$ (3)
+Intuitively, based on the classification predictions $\mathbf{A}$, we partition the action queries decoded by the decoder $\mathbf{D}^{\text{act}}$ into two categories and employ inter-query contrastive learning on them, as shown in Fig.
+2 (right).
+The features that are classified as “repetitive actions” form the positive set $S^{+}$, while the other features form the negative set $S^{-}$.
+Then we apply contrastive learning using InfoNCE loss [12] $\mathcal{L}_{\text{ctrs}}$ over the representation space:
+$\begin{split}\mathcal{L}_{\text{ctrs}}&=-\sum\limits_{i\in S^{+}}{\rm log}\left(\frac{\mathcal{L}_{+}}{\mathcal{L}_{+}+\mathcal{L}_{-}}\right),\\ \mathcal{L}_{+}&=\sum\limits_{s\in S^{+},s\neq i}{\rm exp}(\mathbf{D}^{\text{act}}_{i}\cdot\mathbf{D}^{\text{act}}_{s})/\tau,\\ \mathcal{L}_{-}&=\sum\limits_{s\in S^{-}}{\rm exp}(\mathbf{D}^{\text{act}}_{i}\cdot\mathbf{D}^{\text{act}}_{s})/\tau,\end{split}$ (5)
 where $\tau$ is the temperature parameter, and $\cdot$ denotes inner product.
-3.5 Training Loss
-The overall loss function is defined as:
-$\mathcal{L}=\lambda_{\text{cls}}\mathcal{L}_{\text{cls}}+\lambda_{\text{pos}}\mathcal{L}_{\text{pos}}+\lambda_{\text{ctrs}}\mathcal{L}_{\text{ctrs}},$ (4)
-where $\lambda_{\text{cls}}$, $\lambda_{\text{pos}}$, $\lambda_{\text{ctrs}}$ are the coefficients of each loss term, respectively.
+3.6 Training
+We train our model in an end-to-end manner using the overall loss function:
+$\mathcal{L}=\lambda_{\text{Hungarian}}\mathcal{L}_{\text{Hungarian}}+\lambda_{\text{ctrs}}\mathcal{L}_{\text{ctrs}},$ (6)
+where $\lambda_{\text{Hungarian}},\lambda_{\text{ctrs}}\in\mathbb{R}$ are the coefficients.
+Following DETR [3], we also found it helpful to use auxiliary losses in the decoder during training.
+Specifically, we add $\mathcal{L}_{\text{Hungarian}}$ loss on the predictions from the prediction heads right after the encoder i.e.
+$\tilde{\mathbf{A}}$ and $\tilde{\mathbf{P}}$, and add prediction heads and $\mathcal{L}_{\text{Hungarian}}$ loss after each decoder layer.
+All prediction heads share their parameters.
+This ensures the model focuses on the correct features at each stage consistently, thereby accelerating convergence.
 4 Experiments
@@
 This dataset stands out due to its extensive video lengths, significant variations in the average motion cycle, and a higher number of repetitive cycles compared to prior datasets [9, 33, 18, 47].
-We use the start and end positions of each action instance $\hat{\mathbf{P}}_{q}$ provided by the annotations, and disregard the class action labels.
+We use the start and end positions of each action instance to compute the GT label $\hat{\mathbf{P}}$ provided by the annotations.
 We train our model on the RepCountA train set and select the best model on the validation set.
 We report the evaluation results on the test set.
+Table 1: Comparison to the state-of-the-arts on RepCountA [14] dataset.
+We compare with SOTA action recognition/segmentation methods (top block), TRC methods (second block), and action detection methods (third block with $\dagger$).
+We further report MAE and OBO metrics for short-, medium-, and long-period test actions.
+Backbone ${\rm MAE}\downarrow$ ${\rm OBO}\uparrow$ ${\rm MAE_{s}}\downarrow$ ${\rm OBO_{s}}\uparrow$ ${\rm MAE_{m}}\downarrow$ ${\rm OBO_{m}}\uparrow$ ${\rm MAE_{l}}\downarrow$ ${\rm OBO_{l}}\uparrow$
+X3D [10] X3D 0.9105 0.1059 - - - - - -
+TANet [26] TANet 0.6624 0.0993 - - - - - -
+VideoSwinT [25] ViT 0.5756 0.1324 - - - - - -
+GTRM [15] I3D 0.5267 0.1589 - - - - - -
+RepNet [9] ResNet 0.5865 0.2450 0.7793 0.0930 0.5893 0.1591 0.4549 0.4062
+Zhang et al. [47] 3D-ResNext 0.8786 0.1554 - - - - - -
+TransRAC [14] ViT 0.4891 0.2781 0.5789 0.0233 0.4696 0.2955 0.4420 0.4375
+Li et al. [19] ViT 0.3841 0.3860 - - - - - -
+TadTR [24] $\dagger$ I3D 1.1314 0.0662 0.8364 0.0233 1.1591 0.0000 1.3106 0.1406
+ActionFormer [45] $\dagger$ I3D 0.4990 0.2781 0.4164 0.1628 0.3768 0.3409 0.6385 0.3125
+ReAct [34] $\dagger$ TSN 0.4592 0.3509 0.2805 0.1576 0.3037 0.4318 0.6862 0.3906
+Ours TSN 0.2809 0.4570 0.2411 0.1628 0.1792 0.5455 0.3776 0.5938
+Ours I3D 0.3305 0.4437 0.2421 0.2093 0.2012 0.5227 0.4788 0.5469
+Ours ViT 0.2622 0.5430 0.2257 0.2558 0.2002 0.5909 0.3294 0.7031
 UCFRep dataset [47] is a subset of the UCF101 dataset [36], including fitness videos and daily life videos.
@@
 Metrics.
-We compute two commonly used metrics, OBO and MAE  [13, 9, 46], to evaluate the model performance.
+Following previous works [14, 9, 47, 19], we compute two commonly used metrics, OBO and MAE, to evaluate the model performance.
 OBO (Off-By-One count error) measures the percentage that the predicted count is within the GT count ±$\pm 1$ range.
@@
 Formally,
-${\rm OBO}=\frac{1}{M}\sum_{i=1}^{M}\left|N_{i}-\hat{N}_{i}\leq 1\right|,\quad{\rm MAE}=\frac{1}{M}\sum_{i=1}^{M}\frac{\left|N_{i}-\hat{N}_{i}\right|}{N_{i}},$ (5)
-where ${N}_{i}$ and $\hat{N}_{i}$ are the predicted and GT counts for the $i^{th}$ test video, respectively, and $M$ is the total number of test videos.
+${\rm OBO}$ $=\frac{1}{M}\sum_{m=1}^{M}\left|N_{m}-\hat{N}_{m}\leq 1\right|,$ (7)
+${\rm MAE}$ $=\frac{1}{M}\sum_{m=1}^{M}\frac{\left|N_{m}-\hat{N}_{m}\right|}{N_{m}},$ (8)
+where $M$ is the total number of test videos, ${N}_{m}$ and $\hat{N}_{m}$ are the predicted and GT counts for the $m^{th}$ test video, respectively.
 To better evaluate the performance of different models in recognizing actions with varying periods, we expand the evaluation metrics with three variants for the OBO and MAE metrics.
@@
 4.2 Implementation Details
-We implement our approach with different backbone video feature extractors respectively, including TSN [40], I3D [4], and ViT-B [8].
-For the Transformer architecture, we employ a 2-layer encoder and a 4-layer decoder, both with 8-head attention mechanisms.
+We implement our approach with different backbone video feature extractors respectively, including TSN [41], I3D [4], and ViT [8].
+For encoder-decoder Transformer, we set ${L}^{\text{enc}}=2$ and ${L}^{\text{dec}}=4$, both with 8-head attention mechanisms.
 The feature dimension is set to $C=512$.
@@
 We utilize the AdamW optimizer [27] with a learning rate of $0.002$, a batch size of $64$, and train the model for $80$ epochs.
-We set $\lambda_{cls}=1.0$, $\lambda_{pos}=5.0$, $\lambda_{gIoU}=0.4$, $\lambda_{ctrs}=1.0$, confidence threshold $\alpha=0.2$.
+We set $\lambda_{\text{Hungarian}}=1.0$, $\lambda_{\text{L1}}=5.0$, $\lambda_{\text{gIoU}}=0.4$, $\lambda_{\text{ctrs}}=1.0$, confidence threshold $\alpha=0.2$.
 We provide further implementation details in the supplementary.
 4.3 Comparison to State-of-the-arts
+Table 2: Generalization comparison with SOTA TRC methods on UCFRep [47] dataset. MAE and OBO metrics for short-, medium-, and long-period actions are also reported.
+Backbone ${\rm MAE}\downarrow$ ${\rm OBO}\uparrow$ ${\rm MAE_{s}}\downarrow$ ${\rm OBO_{s}}\uparrow$ ${\rm MAE_{m}}\downarrow$ ${\rm OBO_{m}}\uparrow$ ${\rm MAE_{l}}\downarrow$ ${\rm OBO_{l}}\uparrow$
+RepNet [9] ResNet 0.5336 0.2984 0.6219 0.1739 0.4825 0.3600 0.4996 0.5000
+TransRAC [14] ViT 0.6180 0.3143 0.6296 0.1951 0.5842 0.4250 0.6784 0.4118
+Li et al. [19] 3D-ResNext 0.5227 0.3500 - - - - - -
+Ours TSN 0.6016 0.2959 0.7069 0.0488 0.5777 0.4250 0.4039 0.5882
+Ours I3D 0.5194 0.3980 0.4945 0.2195 0.5865 0.4250 0.4216 0.7647
+Ours ViT 0.5435 0.4184 0.5657 0.1951 0.4625 0.5500 0.6804 0.6471
+Table 3: Comparison of computational complexity. We benchmark all the methods with inference on the RepCountA dataset.
+       $T$        Params (M)        FLOPs (G)
+       RepNet [9]        64        20.47        92.40
+       TransRAC [14]        64        14.43        100.19
+       Ours        64        18.87        1.44
+       Ours        512        18.87        6.06
 Results on RepCountA dataset.
@@
 We change their output layers accordingly and train them on RepCountA.
-Table 1: Comparison to the state-of-the-arts on RepCountA [13] dataset.
-We compare with SOTA action recognition/segmentation methods (top block), TRC methods (second block), and action detection methods (third block with normal-$\dagger$).
-We further report MAE and OBO metrics for short-, medium-, and long-period test actions.
-Backbone ${\rm MAE}\downarrow$ ${\rm OBO}\uparrow$ ${\rm MAE_{s}}\downarrow$ ${\rm OBO_{s}}\uparrow$ ${\rm MAE_{m}}\downarrow$ ${\rm OBO_{m}}\uparrow$ ${\rm MAE_{l}}\downarrow$ ${\rm OBO_{l}}\uparrow$
-X3D [10] X3D 0.9105 0.1059 - - - - - -
-TANet [25] TANet 0.6624 0.0993 - - - - - -
-VideoSwinT [24] Swin-T 0.5756 0.1324 - - - - - -
-GTRM [14] I3D 0.5267 0.1589 - - - - - -
-RepNet [9] ResNet-50 0.5865 0.2450 0.7793 0.0930 0.5893 0.1591 0.4549 0.4062
-Zhang et al. [46] 3D-ResNext101 0.8786 0.1554 - - - - - -
-TransRAC [13] Swin-T 0.4891 0.2781 0.5789 0.0233 0.4696 0.2955 0.4420 0.4375
-Li et al. [18] Swin-T 0.3841 0.3860 - - - - - -
-TadTR [23] $\dagger$ I3D 1.1314 0.0662 0.8364 0.0233 1.1591 0.0000 1.3106 0.1406
-ActionFormer [44] $\dagger$ I3D 0.4990 0.2781 0.4164 0.1628 0.3768 0.3409 0.6385 0.3125
-ReAct [33] $\dagger$ TSN 0.4592 0.3509 0.2805 0.1576 0.3037 0.4318 0.6862 0.3906
-Ours TSN 0.2809 0.4570 0.2411 0.1628 0.1792 0.5455 0.3776 0.5938
-Ours I3D 0.3305 0.4437 0.2421 0.2093 0.2012 0.5227 0.4788 0.5469
-Ours ViT-B 0.2622 0.5430 0.2257 0.2558 0.2002 0.5909 0.3294 0.7031
 As shown in Tab.
@@
 To adapt them to the TRC task, we modify their output layer to incorporate class-agnostic supervision accordingly.
-However, relying solely on class-agnostic supervision does not provide them with the ability to dynamically identify repetitive cycles based on the input videos.
+However, relying solely on class-agnostic supervision does not support them to dynamically identify repetitive cycles based on the input videos.
 As a result, their performance on long, slow action instances is generally inferior compared to dedicated TRC approaches.
@@
 The results demonstrate the effectiveness and generalization capability of our method.
-Table 2: Generalization comparison with SOTA TRC methods on UCFRep [46] dataset. MAE and OBO metrics for short-, medium-, and long-period actions are also reported.
-Backbone ${\rm MAE}\downarrow$ ${\rm OBO}\uparrow$ ${\rm MAE_{s}}\downarrow$ ${\rm OBO_{s}}\uparrow$ ${\rm MAE_{m}}\downarrow$ ${\rm OBO_{m}}\uparrow$ ${\rm MAE_{l}}\downarrow$ ${\rm OBO_{l}}\uparrow$
-RepNet [9] ResNet-50 0.5336 0.2984 0.6219 0.1739 0.4825 0.3600 0.4996 0.5000
-TransRAC [13] Swin-T 0.6180 0.3143 0.6296 0.1951 0.5842 0.4250 0.6784 0.4118
-Li et al. [18] 3D-ResNext101 0.5227 0.3500 - - - - - -
-Ours TSN 0.6016 0.2959 0.7069 0.0488 0.5777 0.4250 0.4039 0.5882
-Ours I3D 0.5194 0.3980 0.4945 0.2195 0.5865 0.4250 0.4216 0.7647
-Ours ViT-B 0.5435 0.4184 0.5657 0.1951 0.4625 0.5500 0.6804 0.6471
+Figure 3: Qualitative results on RepCountA dataset.
+Each colored block represents a GT or predicted action instance.
+TransRAC represents the results by density map, and the final count value is obtained by summing the values in the density map.
+Figure 4: Visualization of failure case on RepCountA [14] dataset.
+Due to the excessive zooming in, the legs of the human body are truncated, making a large difference in the action motion feature, and resulting in several missed cycle counts.
+Each colored block represents a GT or predicted action instance.
+Figure 5: Qualitative results on UCFRep dataset.
+Each colored block represents a GT or predicted action instance.
+TransRAC represents the results by density map, and the final count value is obtained by summing the values in the density map.
+The vertical lines in the labels represent the time points at which the actions begin since only the starting point annotations are provided in UCFRep [47].
+Table 4: Effect of DAQ and ICL modules on RepCountA [14] dataset. (a) ablates DAQ strategy. (b) ablates the ICL strategy. (c) is our full model.
+${\rm MAE}\downarrow$ ${\rm OBO}\uparrow$ ${\rm MAE_{s}}\downarrow$ ${\rm OBO_{s}}\uparrow$ ${\rm MAE_{m}}\downarrow$ ${\rm OBO_{m}}\uparrow$ ${\rm MAE_{l}}\downarrow$ ${\rm OBO_{l}}\uparrow$
+(a) w/o DAQ 0.3542 0.4172 0.2624 0.2093 0.2515 0.5227 0.4864 0.4844
+(b) w/o ICL 0.4035 0.4040 0.2448 0.2093 0.3106 0.4545 0.5740 0.5000
+(c) Ours (full) 0.2809 0.4570 0.2411 0.1628 0.1792 0.5455 0.3776 0.5938
 Efficiency.
@@
 Unfortunately, due to computational limitations, we are unable to obtain the results of RepNet [9] and TransRAC [14] when $T=512$.
-Table 3: Comparison of computational complexity. We benchmark all the methods with inference on the RepCountA dataset.
-$T$ Params (M) FLOPs (G)
-RepNet [9] 64 20.47 92.40
-TransRAC [13] 64 14.43 100.19
-Ours 64 18.87 1.44
-Ours 512 18.87 6.06
 4.4 Qualitative Results
-We visualize the predictions of our approach and baseline method TransRAC [13] on the RepCountA [13] dataset to better demonstrate our performance improvement in Fig.
+We visualize the predictions of our approach and baseline method TransRAC [14] on the RepCountA [14] dataset in Fig.
 3.
@@
 However, this approach suffers from a lack of interpretability and finally results in miscounting.
-In contrast, our approach not only produces an accurate final count but also precisely localizes the action start and end positions (colored blocks) in most cases, aligning with the GT labels.
-It can be seen that our method exhibits robustness to changes in viewpoint, background noise, and sudden interruptions.
-For instance, in the second case of Fig.
-3, we accurately and robustly estimate the time positions even with viewpoint changes as well as a sudden interruption in the middle of the timeline.
+In contrast, our approach not only produces an accurate final count but also correctly localizes the action start and end positions (colored blocks) in most cases.
+In addition, our method exhibits robustness to changes in viewpoint, background noise, and sudden interruptions, e.g., in the second case of Fig.
+3, we accurately estimate the time positions even with viewpoint changes as well as a sudden interruption in the middle of the timeline.
+Please refer to our supplementary video for more qualitative results.
+We present the video result of the first case in Fig.
+3, which shows that while the subject performs pull-ups, there are moments of talking and arm waving.
+Our method selectively counts only the pull-up actions, as they are the primary focus and consistent with the main action instances of the video, excluding dissimilar actions like “waving”, validating the robustness and generalization ability of our method.
 In addition, our query-based representation offers excellent interpretability, making it easy to identify the issues in case of failures.
@@
 Due to the excessive zooming in, the legs of the human body are truncated, making a large difference in the action motion feature, and resulting in several missed cycle counts.
-Figure 3: Qualitative results on RepCountA dataset.
-Each colored block represents a GT or predicted action instance.
-TransRAC represents the results by density map, and the final count value is obtained by summing the values in the density map.
-Figure 4: Visualization of failure case on RepCountA [13] dataset.
-Due to the excessive zooming in, the legs of the human body are truncated, making a large difference in the action motion feature, and resulting in several missed cycle counts.
-Each colored block represents a GT or predicted action instance.
 We further illustrate the generalization performance of the proposed method on the unseen UCFRep [47] test set in Fig.
@@
 We attribute the performance advantage to the proposed DAQ and ICL designs, which empower the model to adaptively adjust the action queries based on the input video features and effectively localize similar (repetitive) action instances, distinguishing them from the background noise actions.
-Please refer to our supplementary video for more qualitative results which are robust to varying cycle lengths, sudden interruptions, and multiple distracting action classes.
-Figure 5: Qualitative results on UCFRep dataset.
-Each colored block represents a GT or predicted action instance.
-TransRAC represents the results by density map, and the final count value is obtained by summing the values in the density map.
-The vertical lines in the labels represent the time points at which the actions begin since only the starting point annotations are provided in UCFRep [46].
 4.5 Ablation Study
@@
 4 presents the results on RepCountA [14] using the TSN backbone.
-In ablation (a), we substitute the proposed DAQ module with a static content query, where the content queries $\mathbf{Q}^{\text{cnt}}$ are also learnable variables, instead of using the content queries filtered out by the encoder and query selection modules.
+In ablation (a), we substitute the proposed DAQ module with a static action query, where the action queries $\mathbf{Q}^{\text{act}}$ are also learnable variables, instead of using the action queries filtered out by the encoder and query selection modules.
 In ablation (b), we eliminate the ICL design among the action queries.
 The results demonstrate that both DAQ and ICL contribute to performance improvement, particularly in terms of counting medium and long actions.
-Table 4: Effect of DAQ and ICL modules on RepCountA [13] dataset. (a) ablates DAQ strategy. (b) ablates the ICL strategy. (c) is our full model.
-${\rm MAE}\downarrow$ ${\rm OBO}\uparrow$ ${\rm MAE_{s}}\downarrow$ ${\rm OBO_{s}}\uparrow$ ${\rm MAE_{m}}\downarrow$ ${\rm OBO_{m}}\uparrow$ ${\rm MAE_{l}}\downarrow$ ${\rm OBO_{l}}\uparrow$
-(a) w/o DAQ 0.3542 0.4172 0.2624 0.2093 0.2515 0.5227 0.4864 0.4844
-(b) w/o ICL 0.4035 0.4040 0.2448 0.2093 0.3106 0.4545 0.5740 0.5000
-(c) Ours (full) 0.2809 0.4570 0.2411 0.1628 0.1792 0.5455 0.3776 0.5938
 Confidence threshold $\alpha$.
-We evaluate the impact of different confidence thresholds for the “action of interest” on the final counting results in Fig.
-6 (green curve).
+We report the impact of different confidence thresholds for classifying “repetitive actions” on the final counts in Fig.
+6.
+We illustrate the MAE (left) and OBO (right) curves for our method (green curve) and the TransRAC [14] approach (blue curve) across various confidence thresholds $\alpha$.
+The metrics for TransRAC are derived by binarizing its output density map and summing the results to obtain the final count.
 The results indicate that setting the threshold within the range of $0.2$ to $0.4$ yields similar performance for our method.
@@
 Figure 6: Results of different confidence thresholds of our method and TransRAC [14].
-We depict the MAE (left) and OBO (right) curves of our method (green curve) and the TransRAC (blue curve) approach with regard to different confidence thresholds.
+We depict the MAE (left) and OBO (right) curves of our method (green curve) and the TransRAC (blue curve) approach concerning different confidence thresholds.
 The metrics of TransRAC are obtained by binarizing the density map of TransRAC output and then summing to obtain the final count.
 5 Conclusion
-In conclusion, our study introduces an innovative perspective for the TRC task, which reduces the computational complexity and maintains robustness across varying action periods and video lengths.
-To tackle the challenge of open-set action categories, we propose DAQ, which improves generalization across different actions.
-To recognize actions with identical content, we propose ICL, which facilitates learning repetitive action representations and distinguishing them from distractions.
-Integrating DAQ and ICL, our method can identify contextually similar action instances that are adaptively conditioned on the video content.
-Experimental results on challenging benchmarks demonstrate the superiority of our approach compared to SOTAs, in terms of both accuracy and efficiency.
-Furthermore, our approach adeptly balances the demands of diverse action speeds and video durations, establishing a solid foundation for practical implementations in real-world scenarios.
+In conclusion, we provide an innovative perspective for tackling TRC tasks, which reduces computational complexity and maintains robustness across varying action lengths.
+To address open-set action categories, we propose DAQ for improved generalization, and ICL for recognizing repetitive actions and distinguishing them from distractions.
+Integrating DAQ and ICL, our method adaptively identifies contextually similar action instances.
+Experimental results on challenging benchmarks demonstrate our approach’s superiority over SOTAs in both accuracy and efficiency, while adeptly balancing diverse action speeds and video durations, establishing a solid foundation for practical implementations in real-world scenarios.
+[…]
+Acknowledgements This work was supported by National Science and Technology Major Project (2022ZD0114904).
+[…]
+Data Availability Statements.
+The RepCountA [14] and UCFRep [47] datasets that support the findings of this study are publicly available on GitHub at https://github.com/SvipRepetitionCounting/TransRAC and https://github.com/Xiaodomgdomg/Deep-Temporal-Repetition-Counting, respectively.

### v4

@@
-
+Efficient Action Counting with Dynamic Queries
 Abstract
 Most existing methods rely on the similarity correlation matrix to characterize the repetitiveness of actions, but their scalability is hindered due to the quadratic computational complexity.
-In this work, we introduce a novel approach that employs an action query representation to localize repeated action cycles with linear computational complexity.
+In this work, we introduce a novel approach that employs an action query representation to localize class-agnostic repeated action cycles with linear computational complexity.
 Based on this representation, we develop two key components to tackle the essential challenges of temporal repetition counting.
-Firstly, to facilitate open-set action counting, we propose the dynamic update scheme on action queries.
-Unlike static action queries, this approach dynamically embeds video features into action queries, offering a more flexible and generalizable representation.
-Secondly, to distinguish between actions of interest and background noise actions, we incorporate inter-query contrastive learning to regularize the video representations corresponding to different action queries.
-As a result, our method significantly outperforms previous works, particularly in terms of long video sequences, unseen actions, and actions at various speeds.
+Firstly, to tackle open-set action counting, we define two action classes: “repetitive actions” and “others”.
+Instead of manually defining the repetitive action class, we propose a dynamic action query strategy.
+Here, each action query directly represents an extracted video feature, allowing the repetitive actions of interest to be dynamically defined based on the video content itself.
+Secondly, to distinguish these repetitive action queries from others, we propose inter-query contrastive learning.
+This performs contrastive clustering over the queries, pulling similar action patterns together while pushing apart those related to background or unrelated movements.
+As a result, queries classified as “repetitive actions” are considered as repetitive cycles, which are then used for counting.
+Thanks to the query-based representation and contrastive learning strategy, our method significantly outperforms previous works on accuracy while being more lightweight and time-efficient.
 On the challenging RepCountA benchmark, we outperform the state-of-the-art method TransRAC by 26.5% in OBO accuracy, with a 22.7% mean error decrease and 94.1% computational burden reduction.
-Code will be publicly available.
+Code and models are publicly available at project page.
 keywords:
@@
 Temporal periodicity is a ubiquitous phenomenon in the natural world.
-Temporal Repetition Counting (TRC) aims to accurately measure the number of repetitive action cycles within a given video and holds significant potential for applications such as fitness monitoring [11] and motion generation [49].
+Temporal Repetition Counting (TRC) aims to accurately measure the number of repetitive action cycles within a given video and holds significant potential for applications such as fitness monitoring [fieraru2021aifit] and motion generation [zhu2023human].
 Figure 1: Conceptual workflow comparison of (a) similarity-matrix-based methods and (b) our proposed action query-based method.
@@
 This novel formulation reduces the complexity from quadratic to linear as $O(TC)$.
-Pioneer methods  [17, 1, 7, 39, 29, 38, 6] represent time-series video data as one-dimensional signals and employ spectral analysis techniques such as the Fourier transform.
+Pioneer methods  [laptev2005periodic, azy2008segmentation, cutler2000robust, tsai1994cyclic, pogalin2008visual, thangali2005periodic, chetverikov2006motion] represent time-series video data as one-dimensional signals and employ spectral analysis techniques such as the Fourier transform.
 While suitable for short videos with fixed periodic cycle lengths, these methods struggle to handle real-world scenarios with varying cycle lengths and sudden interruptions.
-Recent studies shift to deep learning-based methods [18, 9, 14, 47, 19] and show promising performances.
-Notably, most of these methods, such as RepNet [9] and TransRAC [14], utilize a temporal similarity correlation matrix to depict repetitiveness, as illustrated in Fig.
-1 (a).
+Recent studies shift to deep learning-based methods [levy2015live, dwibedi2020counting, hu2022transrac, zhang2020context, li2024repetitive] and show promising performances.
+Notably, most of these methods, such as RepNet [dwibedi2020counting] and TransRAC [hu2022transrac], utilize a temporal similarity correlation matrix to depict repetitiveness, as illustrated in Fig.˜1 (a).
 Nevertheless, the computational complexity of this representation grows quadratically with the number of input frames $T$, highlighting a significant gap in scalability that hinders their application to real-world scenarios of varying action periods and dynamics.
-Recent progress in action detection [34, 24] introduces an efficient representation of action periods by associating each action instance with an action query, similar to DETR [3].
-Inspired by this, we propose to formulate the TRC problem as a set prediction task where the goal is to detect every action cycle by representing it as an action query, as illustrated in Fig.
-1 (b).
-Based on this query representation, we use a Transformer encoder-decoder network [50] to detect repetitive action instances and their temporal positions, defined by their midpoints and durations.
-This novel formulation reduces the complexity from quadratic to linear  [^1] In implementation, we employ deformable attention modules as proposed in DeformableDETR [50].
+Recent progress in action detection [shi2022react, liu2022end] introduces an efficient representation of action periods by associating each action instance with an action query, similar to DETR [carion2020end].
+Inspired by this, we propose to formulate the TRC problem as a set prediction task where the goal is to detect every action cycle by representing it as an action query, as illustrated in Fig.˜1 (b).
+Based on this query representation, we use a Transformer encoder-decoder network [zhu2020deformable] to detect repetitive action instances and their temporal positions, defined by their midpoints and durations.
+This novel formulation reduces the complexity from quadratic to linear  [^1] In implementation, we employ deformable attention modules as proposed in DeformableDETR [zhu2020deformable].
 and enables counting long videos with varying action periods.
-However, directly applying the action detection approaches [24, 45, 34] to the TRC problem proves inadequate (Tab.
-1) in addressing two distinctive challenges unique to TRC.
+However, directly applying the action detection approaches [liu2022end, zhang2022actionformer, shi2022react] to the TRC problem proves inadequate (Tab.˜1) in addressing two distinctive challenges unique to TRC.
 These challenges underscore the complexity of TRC, highlighting why TRC is not merely another action detection task but requires a nuanced approach that considers the unique nature of repetitive actions.
@@
     TRC requires recognizing action instances with identical content, while detection does not.
-As a result, approaching TRC as a simple action detection task results in inferior performance as shown in Sec.
-4.3.
+As a result, approaching TRC as a simple action detection task results in inferior performance as shown in Sec.˜4.3.
 In contrast, we propose two novel strategies to address these challenges and redefine the framework for TRC.
-In response to the first challenge, we propose Dynamic Action Query (DAQ), which adaptively updates the action query using content features extracted from the video (Fig.
-2).
-This mechanism allows the decoder to attend to the repetitive actions based on the input video contents in a dynamic and contextually aware manner, without the need for manual definition.
-To tackle the second challenge, we define two action classes: “repetitive actions”, which are dynamically defined by DAQ, and the “others ($\emptyset$)” class.
-We further propose Inter-query Contrastive Learning (ICL) which ensures that primary repetitive action cycles are grouped together in the learned representation space while being separated from other distractors, such as background noise.
-The integration of two core components (DAQ and ICL) ensures that the action instances are identified adaptively based on the video content and their contextual similarity.
+In response to the first open-set challenge, we define two action classes: “repetitive actions” and “others ($\emptyset$)”.
+Instead of manually defining the repetitive action class, we propose a Dynamic Action Query (DAQ) strategy, which adaptively updates the action query using content features extracted from the video (Fig.˜2) This mechanism allows the decoder to attend to the repetitive actions based on the input video contents in a dynamic, contextually aware and class-agnostic manner.
+To tackle the second challenge, we further propose Inter-query Contrastive Learning (ICL), which ensures that primary repetitive action cycles of interest are grouped together in the learned representation space while being separated from other distractors, such as background noise.
+The integration of two core components (DAQ and ICL) ensures that the action instances are identified adaptively based on the video content and their contextual similarity, making our approach class-agnostic.
 In other words, our queries are designed to localize the contextually similar action instances, which aligns exactly with the definition of repetition counting.
@@
   1. 1.
-    We provide a novel perspective to tackle the TRC problem using a simple yet effective representation for action cycles.
-Our approach reduces the computational complexity from quadratic to linear and is robust to varying action periods and video lengths.
+    We provide a novel perspective to tackle the TRC problem using a simple yet effective representation for action cycles, i.e.
+action query.
+Our approach reduces the computational complexity from quadratic to linear and is more lightweight and time-efficient.
   2. 2.
@@
   4. 4.
-    By addressing the scalability issue and unique challenges of TRC, our method on two challenging benchmarks notably surpasses state-of-the-art (SOTA) methods in terms of both accuracy and efficiency.
+    Our method notably surpasses state-of-the-art (SOTA) methods in terms of both accuracy and efficiency on two challenging benchmarks.
 Notably, our method strikes an effective balance in handling various action periods and video lengths, offering a significant leap forward in the practical application of TRC technologies.
@@
 2.1 Temporal Repetition Counting
-Traditional methods [17, 1, 7, 39, 29, 38, 6] frequently employ spectral or frequency domain techniques for the analysis of repetitive sequences, thereby preserving the underlying repetitive motion structures.
+Traditional methods [laptev2005periodic, azy2008segmentation, cutler2000robust, tsai1994cyclic, pogalin2008visual, thangali2005periodic, chetverikov2006motion] frequently employ spectral or frequency domain techniques for the analysis of repetitive sequences, thereby preserving the underlying repetitive motion structures.
 While these conventional approaches are capable of effectively handling simple motion sequences or those characterized by fixed periodicity, they prove inadequate when confronted with non-stationary motion sequences encountered in real-world scenarios.
-In contrast, deep-learning-based approaches [18, 9, 14, 47, 19] have demonstrated remarkable performance improvements.
-Notably, RepNet [9] and TransRAC [14] leverage temporal similarity matrices of actions to construct models for counting temporal repetitions.
+In contrast, deep-learning-based approaches [levy2015live, dwibedi2020counting, hu2022transrac, zhang2020context, li2024repetitive] have demonstrated remarkable performance improvements.
+Notably, RepNet [dwibedi2020counting] and TransRAC [hu2022transrac] leverage temporal similarity matrices of actions to construct models for counting temporal repetitions.
 However, these similarity-matrix-based methods are not scalable for long videos due to their quadratic computational complexity.
-Another research line involves predicting the start and end points of each cycle [47] from coarse to fine.
+Another research line involves predicting the start and end points of each cycle [zhang2020context] from coarse to fine.
 Nevertheless, its practicality is hindered by the requirement for over 30 forward passes to count iteratively from a single video.
@@
 2.2 Temporal Action Detection
-The field of temporal action detection [31, 48, 5, 45, 22] is typically classified into two categories: anchor-based methods, and anchor-free methods.
-Anchor-based methods [44, 20, 30] generate multiple anchors, subsequently classifying these anchors to determine the action boundaries.
-Anchor-free methods  [2, 35, 43, 21] predict action instances by directly regressing the boundary and the center point of an action instance.
-With the rapid development of Transformer technology, DETR [3] is introduced for object detection task [50, 28, 23, 46] and gains increasing popularity with promising performance.
-This paradigm promotes the study in many fields such as the action detection tasks [24, 37, 40, 42].
+The field of temporal action detection [redmon2016you, zhao2017temporal, chao2018rethinking, zhang2022actionformer, lin2019bmn] is typically classified into two categories: anchor-based methods, and anchor-free methods.
+Anchor-based methods [zeng2019graph, li2021three, qing2021temporal] generate multiple anchors, subsequently classifying these anchors to determine the action boundaries.
+Anchor-free methods  [buch2019end, shou2017cdc, yuan2017temporal, lin2021learning] predict action instances by directly regressing the boundary and the center point of an action instance.
+With the rapid development of Transformer technology, DETR [carion2020end] is introduced for object detection task [zhu2020deformable, meng2021conditional, liu2022dab, zhang2022dino] and gains increasing popularity with promising performance.
+This paradigm promotes the study in many fields such as the action detection tasks [liu2022end, tan2021relaxed, vaswani2017attention, wang2021oadtr].
 These methods establish a direct connection between action queries and the predicted action instances, enabling them to accurately predict the temporal boundaries of actions.
@@
 3.1 Preliminary
-DETR [3] is a pioneering object detection framework that builds upon the Transformer encoder-decoder architecture [40].
-The overall DETR architecture [3] consists of three main components: a backbone to extract image features, an encoder-decoder Transformer, and the detection heads, i.e.
+DETR [carion2020end] is a pioneering object detection framework that builds upon the Transformer encoder-decoder architecture [vaswani2017attention].
+The overall DETR architecture [carion2020end] consists of three main components: a backbone to extract image features, an encoder-decoder Transformer, and the detection heads, i.e.
 feed-forward network (FFN) that makes the final detection prediction.
@@
 Given the input image feature maps extracted by a CNN backbone, e.g.
-ResNet [13], DETR exploits a standard Transformer encoder-decoder network to transform the feature maps to be features of a set of object queries.
+ResNet [he2016deep], DETR exploits a standard Transformer encoder-decoder network to transform the feature maps to be features of a set of object queries.
 An FFN and a linear projection are added on top of the object query features as the detection heads.
@@
 Figure 2: Overview and detailed architecture design of our method.
-(Left) We utilize a DETR-inspired framework to predict the number of repetitive action cycles in an RGB video sequence.
-The architecture consists of a backbone network $\Phi(\cdot)$ for feature extraction, an encoder-decoder Transformer for processing query features, and prediction heads for action classification.
-By counting the queries classified as “repetitive actions”, we get the total count $N$.
-(Right) The detailed workflow within the black box on the left.
-Given the video features $\mathbf{F}$ as input, the encoder $\mathcal{E}(\cdot)$ produces two sets of query features, i.e.
-action queries and position queries.
-A query selection module then screens these queries, retaining only the most relevant ones for forwarding to the decoder $\mathcal{D}(\cdot)$.
-This process utilizes action classification results from prediction heads positioned after the encoder.
-The decoder adopts the Dynamic Action Query (DAQ) strategy, utilizing the selected action queries as priors, while incorporating a learnable position query and attending to the encoder’s output.
+(Left) We utilize a DETR-inspired framework to predict the number of repetitive action cycles given a video input.
+The framework consists of a backbone network $\Phi(\cdot)$ for feature extraction and the counting module (black box).
+The counting module, including an encoder-decoder Transformer for processing query features and prediction heads for action classification, outputs queries classified as “repetitive actions”, which are then counted to yield the total count NN.
+(Right) The detailed workflow of the counting module.
+Given the video features $\mathbf{F}$ as input, the encoder $\mathcal{E}(\cdot)$ produces a set of query features.
+A query selection module then screens these, retaining only the most relevant ones for forwarding to the decoder $\mathcal{D}(\cdot)$.
+The decoder $\mathcal{D}(\cdot)$ adopts the Dynamic Action Query (DAQ) strategy, using the selected action queries for initialization to dynamically define the “repetitive actions” class.
 The output embeddings from the decoder are then passed to their corresponding prediction head, which estimates the class label $\mathbf{A}$ and temporal location $\mathbf{P}$.
-During training, we employ bipartite matching to uniquely pair each prediction with a GT action cycle $\hat{\mathbf{P}}$, considering both class labels and temporal locations.
-Predictions that fail to match should yield $\emptyset$ class, signifying that they do not represent repetitive actions.
-The DAQ strategy addresses the open-set problem by contextually defining “repetitive actions” based on video features.
-To further distinguish these actions from $\emptyset$ distractions, we propose Inter-query Contrastive Learning (ICL).
-By integrating DAQ and ICL, our method effectively identifies contextually similar actions and excludes distractions, making it well-suited for tackling the challenges of the TRC problem.
+During training, we employ bipartite matching to uniquely pair each prediction with a GT action cycle $\hat{\mathbf{P}}$.
+Predictions that fail to match are assigned the $\emptyset$ class.
+To further distinguish the “repetitive actions” from $\emptyset$ actions, we propose Inter-query Contrastive Learning (ICL).
+Finally, the queries classified as repetitive actions contribute to the total count NN.
 3.2 Overview
-Given an RGB video sequence with $T$ frames, the TRC task aims to predict an integer $N$ indicating the number of detected repetitive action cycles.
-Drawing inspiration from DETR [3], we streamline the problem as a set detection task in the temporal domain.
-The overall framework consists of three main components: a backbone network $\Phi(\cdot)$, an encoder-decoder Transformer, and prediction heads, as shown in Fig.
-2.
-The backbone extracts video features $\mathbf{F}$, which is transformed to features of a set of action queries by an encoder $\mathcal{E}(\cdot)$ and a decoder $\mathcal{D}(\cdot)$.
-Two prediction heads are added on top of the query features.
-The prediction heads consist of an action head and a position head.
-The action head classifies the query, while the position head estimates its temporal location.
-Building on previous enhancements in DETR models [23], we introduce two distinct types of queries to enhance the query-to-feature similarity and accelerate convergence: action queries and position queries.
-Action queries are designed to capture features pertinent to action classification, whereas position queries concentrate on the temporal dimensions of an action.
-As depicted in the detailed diagram on the right side of Fig.
-2, two specialized prediction heads (i.e.
-action head and position head) are tasked with processing these respective sets of queries to generate predictions.
-To further accelerate convergence, we introduce a post-encoder query selection module that acts as an initial sieve, filtering out the majority of queries unlikely to represent repetitive actions before forwarding the most promising ones to the decoder.
-This enhances the efficiency and focus of the subsequent prediction process.
-Considering the two challenges of the TRC tasks discussed in Sec.
-1, we propose Dynamic Action Query (DAQ) strategy to address the open-set problem.
-DAQ directly uses dynamically updated action query features from the encoder to initialize the action query of the decoder.
-This strategy defines what “repetitive actions” should be detected based on the video content features in a dynamic and contextually aware manner.
-To tackle the second challenge, we define two action classes: “repetitive actions” and the “others ($\emptyset$)” class.
-We further propose Inter-query Contrastive Learning (ICL) to cluster repetitive action queries into a positive action set ($S^{+}$) and group the other queries to a negative set ($S^{-}$) in the feature space.
-Integrating DAQ and ICL allows our method to identify contextually similar action instances that are adaptively based on the video content, and exclude other distracting actions at the same time.
-Similar to DETR [3], during training, we employ bipartite matching which uniquely assigns a prediction to a GT action cycle.
+Given an RGB video sequence with $T$ frames, the TRC task aims to predict an integer $N$ indicating the number of detected primary repetitive action cycles, whose class is not predefined.
+Drawing inspiration from DETR [carion2020end], we streamline the problem as an open-set detection task in the temporal domain and propose to use the action query to represent each potential repetitive cycle.
+The overall framework consists of two main components: a backbone network $\Phi(\cdot)$ and the counting module.
+The backbone extracts video features $\mathbf{F}$.
+The counting module, depicted as the black box in Fig.˜2, is composed of a Transformer encoder $\mathcal{E}(\cdot)$, a decoder $\mathcal{D}(\cdot)$, and prediction heads.
+The extracted $\mathbf{F}$ are transformed into a set of action queries by the Transformer encoder and decoder.
+Two prediction heads are then added on top of these query features, aiming to classify each query and estimate its temporal location.
+Once an action query is classified as a repetitive cycle, the action count $N$ increases by 1.
+Considering the two challenges of the TRC tasks discussed in Sec.˜1, we propose Dynamic Action Query (DAQ) strategy to address the open-set problem.
+Specifically, we define two action classes: “repetitive actions” and the “others ($\emptyset$)” class.
+Instead of manually defining the specific class label for the “repetitive actions”, DAQ directly leverages dynamically updated video features from the encoder to initialize the action query for the decoder.
+This enables the adaptive definition of “repetitive actions” based on the video content, in a contextually aware and class-agnostic manner.
+To tackle the second challenge of distinguishing identical repetitive actions, we further propose Inter-query Contrastive Learning (ICL).
+This approach clusters queries representing identical repetitive actions into a positive action set ($S^{+}$) and groups the other queries into a negative set ($S^{-}$) in the feature space.
+Integrating DAQ and ICL allows our method to identify class-agnostic similar action instances that are adaptively based on the video content, and exclude other distracting actions at the same time.
+Similar to DETR [carion2020end], during training, we employ bipartite matching which uniquely assigns a prediction to a GT action cycle.
 The matching procedure takes into account both the class prediction and the similarity of predicted and GT temporal locations.
@@
 By counting the queries classified as “repetitive actions”, we get the final total count value $N$.
-In the following, we will introduce the network architecture design in Sec. 3.3, the DAQ and ICL modules in Sec. 3.4 and Sec. 3.5, and the model training in Sec. 3.6.
+In the following, we will introduce the network architecture design in Sec.˜3.3, the DAQ and ICL modules in Sec.˜3.4 and Sec.˜3.5, and the model training in Sec.˜3.6.
 3.3 Model Architecture
+As illustrated in Fig.˜2, our model can be divided into two main components: a backbone $\Phi(\cdot)$ and the counting module (black box).
+The backbone extracts features $\mathbf{F}$ from the raw video input.
+The counting module then takes these video features $\mathbf{F}$ as input and ultimately outputs the count N.
+Specifically, the counting module includes a Transformer encoder-decoder, a query selection module, prediction heads, and a bipartite matching module.
+Next, we introduce each of them in detail.
 Backbone.
@@
 Encoder.
-The encoder $\mathcal{E}(\cdot)$ is a classical Transformer [40] architecture which has ${L}^{\text{enc}}$ standard encoder layers.
-The encoder transforms the video features $\mathbf{F}$ into two distinct query types: action queries $\tilde{\mathbf{E}}^{\text{act}}\in\mathbb{R}^{T\times C}$ and position queries $\tilde{\mathbf{E}}^{\text{pos}}\in\mathbb{R}^{T\times C}$, as shown in Fig.
-2 (right).
-This design enhances the query-to-feature similarity and accelerates convergence, where action queries capture features pertinent to action classification.
-On the other hand, position queries concentrate on the temporal dimensions of an action.
-We defer to the supplementary material the detailed encoder architecture.
+The encoder $\mathcal{E}(\cdot)$ is a classical Transformer [vaswani2017attention] architecture which has ${L}^{\text{enc}}$ standard encoder layers.
+The encoder transforms the video features $\mathbf{F}$ into two distinct query types: action queries $\tilde{\mathbf{E}}^{\text{act}}\in\mathbb{R}^{T\times C}$ and position queries $\tilde{\mathbf{E}}^{\text{pos}}\in\mathbb{R}^{T\times C}$, as shown in Fig.˜2 (right).
+Action queries capture features pertinent to action classification, while position queries concentrate on the temporal dimensions of an action.
+These features prepare the decoder for action query initialization, from which it will identify repetitive action cycles.
+Please refer to the supplementary material for the detailed encoder architecture.
 Query selection.
-We add a query selection module before passing the encoder output query features to the decoder.
-We route the query features outputted by the encoder to two respective prediction heads (described later in this section), which independently decode them into predictions for action class and temporal location, as depicted on the right side of Fig.
-2.
+We add a query selection module before passing the encoder output query features to the decoder, as depicted on the right side of Fig.˜2.
+This module aims to perform an initial filtering over the encoded video features $\mathbf{F}$.
+Since the encoder $\mathcal{E}(\cdot)$ produces a large number of tokens $\tilde{\mathbf{E}}^{\text{act}}\in\mathbb{R}^{T\times C}$ and $\tilde{\mathbf{E}}^{\text{pos}}\in\mathbb{R}^{T\times C}$, many of them may correspond to background or distracting content.
+Therefore the query selection module selects a subset of $Q$ informative features from the original $T$ tokens as $\mathbf{E}^{\text{act}}\in\mathbb{R}^{Q\times C}$ and $\mathbf{E}^{\text{pos}}\in\mathbb{R}^{Q\times C}$.
+In practice, we route the query features outputted by the encoder to two respective prediction heads (described later in this section), which independently decode them into predictions for action class and temporal location.
 The action head processes the action queries $\tilde{\mathbf{E}}^{\text{act}}$ to estimate the action class $\tilde{\mathbf{A}}$ for each query, while the position head decodes the position queries $\tilde{\mathbf{E}}^{\text{pos}}$ to their temporal locations $\tilde{\mathbf{P}}$.
@@
 Decoder.
-The decoder $\mathcal{D}(\cdot)$ is also a classical Transformer [40] architecture, consisting of ${L}^{\text{dec}}$ standard decoder layers.
+The decoder $\mathcal{D}(\cdot)$ is also a classical Transformer [vaswani2017attention] architecture, consisting of ${L}^{\text{dec}}$ standard decoder layers.
 It processes a set of action queries $\mathbf{Q}^{\text{act}}\in\mathbb{R}^{Q\times C}$ and position queries $\mathbf{Q}^{\text{pos}}\in\mathbb{R}^{Q\times C}$, while simultaneously attending to the direct output (i.e.
-$\tilde{\mathbf{E}}^{\text{act}}$ and $\tilde{\mathbf{E}}^{\text{pos}}$) from the encoder, as depicted in Fig.
-2 (right).
-Following the practice in DETR [3], these inputs are decoded in parallel across each decoder layer and transformed into the corresponding query features $\mathbf{D}^{\text{act}}\in\mathbb{R}^{Q\times C}$ and $\mathbf{D}^{\text{pos}}\in\mathbb{R}^{Q\times C}$.
+$\tilde{\mathbf{E}}^{\text{act}}$ and $\tilde{\mathbf{E}}^{\text{pos}}$) from the encoder, as depicted in Fig.˜2 (right).
+Following the practice in DETR [carion2020end], these inputs are decoded in parallel across each decoder layer and transformed into the corresponding query features $\mathbf{D}^{\text{act}}\in\mathbb{R}^{Q\times C}$ and $\mathbf{D}^{\text{pos}}\in\mathbb{R}^{Q\times C}$.
 The input position queries $\mathbf{Q}^{\text{pos}}$ are initialized as learnable parameters.
@@
 $\mathbf{Q}^{\text{act}}=\mathbf{E}^{\text{act}}$.
-The details of the DAQ strategy will be elaborated in Sec.
-3.4.
+The details of the DAQ strategy will be elaborated in Sec.˜3.4.
 Using self- and encoder-decoder attention over these embeddings, our method globally reasons selected video frame features together while being able to use the whole frame-wise video feature as context.
@@
 Prediction heads.
-The prediction heads are designed to make the final predictions using the features processed by the encoder-decoder.
+The prediction heads are designed to make the final predictions using the features processed by the encoder-decoder, as shown in Fig.˜2 (right).
 They comprise two networks: an action head and a position head, both of which are Multi-Layer Perceptrons (MLPs).
@@
 Bipartite matching.
-Following the practice of DETR [3], our method produces a fixed-size set of $Q$ predictions in a single pass, where $Q$ is set to be much larger than the typical number of action cycles in a $T$-frame video sequence.
-Note that the $Q$ predictions are in a random order.
-One of the main difficulties of training is to match the predicted action cycles (classes and temporal locations) to the GT action cycles, therefore, optimizing action cycle-specific losses, e.g.
-classification loss, temporal position loss, etc.
-Let us denote GT repetitive action cycle labels by $\hat{\mathbf{Y}}=(\hat{\mathbf{A}},\hat{\mathbf{P}})$.
-$\hat{\mathbf{A}}=\mathbf{1}\in\mathbb{R}^{\hat{N}\times 1}$ denotes the $\hat{N}$ repetitive action class labels, where we assign the value 1 to the “repetitive actions”.
-$\hat{\mathbf{P}}\in\mathbb{R}^{\hat{N}\times 2}$ denotes the GT temporal positions of the $\hat{N}$ repetitive action, represented by $\hat{N}$ sets of midpoints and duration.
-Similarly, we denote the predicted repetitive actions by $\mathbf{Y}=(\mathbf{A},\mathbf{P})$, a set of size $Q$ predictions.
-We set $Q$ much larger than $\hat{N}$ empirically.
-We consider $\hat{\mathbf{Y}}$ also a set of size $Q$ padded with “others ($\emptyset$)”, i.e.
-we assign the value 0 to denote $\emptyset$ class.
-To find a bipartite matching between these two sets $\hat{\mathbf{Y}}$ and $\mathbf{Y}$, we search for a permutation of $Q$ elements $\sigma\in\mathfrak{S}_{Q}$ with the lowest cost:
+Note that our method produces a fixed-size set of $Q$ predictions in a random order, where $Q$ is set to be much larger than the typical number of action cycles in a $T$-frame video sequence.
+To enforce supervision, we employ a matching strategy to pair the predicted action cycles (classes and temporal locations) with the GT action cycles, as illustrated in the bottom right of Fig.˜2.
+[…]
 $\hat{\sigma}=\arg\min_{\sigma\in\mathfrak{S}_{Q}}\sum_{i=1}^{Q}\mathcal{L}_{\text{match}}(\hat{\mathbf{Y}}_{i},\mathbf{Y}_{\sigma(i)}),$ (1)
@@
 where $\mathbbm{1}$ is an indicator function.
-To measure the similarity of the predicted temporal locations and the GT temporal location, we define $\mathcal{L}_{\text{pos}}$ using the linear combination of the $L_{1}$ distance and the generalized Intersection over Union (IoU) loss [32].
+To measure the similarity of the predicted temporal locations and the GT temporal location, we define $\mathcal{L}_{\text{pos}}$ using the linear combination of the $L_{1}$ distance and the generalized Intersection over Union (IoU) loss [rezatofighi2019generalized].
 Overall, the position loss is defined as
@@
 where $\lambda_{\text{gIoU}},\lambda_{\text{L1}}\in\mathbb{R}$ are hyperparamters.
-By employing the Hungarian matching algorithm [16] to optimize Eq.
-1, we can achieve the final optimal matching $\hat{\sigma}$ which uniquely assigns a prediction to a GT action cycle, finding one-to-one matching without duplicates.
+By employing the Hungarian matching algorithm [kuhn1955hungarian] to optimize Eq.˜1, we can achieve the final optimal matching $\hat{\sigma}$ which uniquely assigns a prediction to a GT action cycle, finding one-to-one matching without duplicates.
 Notice that the matching cost between a repetitive action instance and $\emptyset$ doesn’t depend on the prediction, which means that in that case the cost is a constant.
@@
 We define a Hungarian loss for all pairs matched in the previous step to supervise both classification and temporal location predictions, i.e.
-a linear combination of a negative log-likelihood loss for class prediction and a temporal position loss defined in Eq.
-3:
+a linear combination of a negative log-likelihood loss for class prediction and a temporal position loss defined in Eq.˜3:
 $\begin{split}\mathcal{L}_{\text{Hungarian}}(\hat{\mathbf{Y}},\mathbf{Y})&=\sum\limits_{i=1}^{Q}\left[-\text{log}\,{p_{\hat{\sigma}(i)}}{(\hat{\mathbf{A}}_{i})}\right.\\ &+\left.\mathbbm{1}_{\{\hat{\mathbf{A}}_{i}\neq\emptyset\}}\mathcal{L}_{\text{pos}}(\hat{\mathbf{P}}_{i},\mathbf{P}_{\hat{\sigma}(i)})\right],\end{split}$ (4)
-where $\hat{\sigma}$ is the optimal assignment computed in the previous bipartite matching step, i.e. Eq. 1.
+where $\hat{\sigma}$ is the optimal assignment computed in the previous bipartite matching step, i.e. Eq.˜1.
 3.4 Dynamic Action Query
-As discussed in Sec.
-1, the TRC problem requires recognizing open-set action instances depending on the video content, where the action category is not predefined.
-Therefore, we propose the Dynamic Action Query strategy, which adaptively updates the action query $\mathbf{Q}^{\text{act}}$ using the selected encoder action query features $\mathbf{E}^{\text{act}}$, i.e.
-$\mathbf{Q}^{\text{act}}=\mathbf{E}^{\text{act}}$, as shown in Fig.
-2.
-The DAQ strategy, simple yet effective, provides priors for the decoder to dynamically and contextually attend to “repetitive actions” based on the input video content, eliminating the need for manually defining action categories and thereby enhancing the model’s generalization capability.
-This mechanism ensures that the cross-attention module in the decoder pools features from the entire video while focusing on relevant local repetitive features.
-Such dynamic adjustment allows the model to concentrate on pertinent “repetitive actions”, offering a deeper understanding of action queries and improving training convergence.
-In summary, the DAQ approach not only circumvents the need for manual action category definitions but also leverages dynamic, contextually aware priors to enhance both the focus and generalization capabilities of the model.
-Our supplementary material further explores various methods for initializing the queries in the decoder $\mathcal{D}(\cdot)$, confirming that the DAQ strategy is the most effective.
+As discussed in Sec.˜1, the TRC problem requires recognizing open-set action instances depending on the video content, where the action category is not predefined.
+Therefore, we formulate the problem as a binary classification task with two categories: “repetitive actions” and “others ($\emptyset$)”.
+The key idea is that the class of “repetitive actions” is not fixed or manually defined.
+Rather, it is dynamically inferred based on the video content.
+To achieve this, we propose the Dynamic Action Query strategy.
+As shown in Fig.˜2 (top right), DAQ adaptively updates the action query $\mathbf{Q}^{\text{act}}$ by directly assigning it the selected encoder action query features $\mathbf{E}^{\text{act}}$, i.e.
+$\mathbf{Q}^{\text{act}}\leftarrow\mathbf{E}^{\text{act}}$.
+These decoder action queries $\mathbf{Q}^{\text{act}}$ embed the video content, enabling the decoder to dynamically define “repetitive actions” based on the input video content.
+The DAQ strategy, simple yet effective, eliminates the need for manually defining action categories and thereby enhances the model’s generalization capability.
+Besides, we also explore several different methods for initializing the queries in the decoder $\mathcal{D}(\cdot)$ in the supplementary, confirming that the DAQ strategy is the most effective.
 3.5 Inter-query Contrastive Learning
@@
 This requires that the actions of interest we classify exhibit similarity in their motion patterns, while other action queries should have dissimilar representations.
-To tackle this challenge, we propose Inter-query Contrastive Learning to distinguish the action queries.
-Intuitively, based on the classification predictions $\mathbf{A}$, we partition the action queries decoded by the decoder $\mathbf{D}^{\text{act}}$ into two categories and employ inter-query contrastive learning on them, as shown in Fig.
-2 (right).
-The features that are classified as “repetitive actions” form the positive set $S^{+}$, while the other features form the negative set $S^{-}$.
-Then we apply contrastive learning using InfoNCE loss [12] $\mathcal{L}_{\text{ctrs}}$ over the representation space:
+To tackle this challenge, we propose Inter-query Contrastive Learning to distinguish the action queries, as shown in Fig.˜2 (right).
+ICL encourages the model to pull together queries corresponding to the same repetitive actions in the feature space, while pushing apart those related to background or unrelated movements.
+Therefore, we partition the action queries decoded by the decoder $\mathbf{D}^{\text{act}}$ into two categories and employ contrastive learning on them based on the classification predictions $\mathbf{A}$.
+Specifically, features classified as “repetitive actions” form the positive set $S^{+}$, while the other features form the negative set $S^{-}$.
+Then we apply contrastive learning using InfoNCE loss [he2020momentum] $\mathcal{L}_{\text{ctrs}}$ over the representation space:
 $\begin{split}\mathcal{L}_{\text{ctrs}}&=-\sum\limits_{i\in S^{+}}{\rm log}\left(\frac{\mathcal{L}_{+}}{\mathcal{L}_{+}+\mathcal{L}_{-}}\right),\\ \mathcal{L}_{+}&=\sum\limits_{s\in S^{+},s\neq i}{\rm exp}(\mathbf{D}^{\text{act}}_{i}\cdot\mathbf{D}^{\text{act}}_{s})/\tau,\\ \mathcal{L}_{-}&=\sum\limits_{s\in S^{-}}{\rm exp}(\mathbf{D}^{\text{act}}_{i}\cdot\mathbf{D}^{\text{act}}_{s})/\tau,\end{split}$ (5)
@@
 where $\lambda_{\text{Hungarian}},\lambda_{\text{ctrs}}\in\mathbb{R}$ are the coefficients.
-Following DETR [3], we also found it helpful to use auxiliary losses in the decoder during training.
+Following DETR [carion2020end], we also found it helpful to use auxiliary losses in the decoder during training.
 Specifically, we add $\mathcal{L}_{\text{Hungarian}}$ loss on the predictions from the prediction heads right after the encoder i.e.
@@
 4.1 Datasets and Metrics
-RepCountA dataset [14] is currently the largest and most challenging benchmark for the video TRC task  [^3] The RepCountB [14] test subset is proprietary and not publicly available..
+RepCountA dataset [hu2022transrac] is currently the largest and most challenging benchmark for the video TRC task  [^3] The RepCountB [hu2022transrac] test subset is proprietary and not publicly available..
 It is primarily compiled from fitness videos on YouTube, including a wide range of fitness activities conducted in diverse settings, including homes, gyms, and outdoor environments.
-This dataset stands out due to its extensive video lengths, significant variations in the average motion cycle, and a higher number of repetitive cycles compared to prior datasets [9, 33, 18, 47].
+This dataset stands out due to its extensive video lengths, significant variations in the average motion cycle, and a higher number of repetitive cycles compared to prior datasets [dwibedi2020counting, runia2018real, levy2015live, zhang2020context].
 We use the start and end positions of each action instance to compute the GT label $\hat{\mathbf{P}}$ provided by the annotations.
@@
 We report the evaluation results on the test set.
-Table 1: Comparison to the state-of-the-arts on RepCountA [14] dataset.
+Table 1: Comparison to the state-of-the-arts on RepCountA [hu2022transrac] dataset.
 We compare with SOTA action recognition/segmentation methods (top block), TRC methods (second block), and action detection methods (third block with $\dagger$).
@@
   Backbone $\rm MAE$$\downarrow$ $\rm OBO$$\uparrow$ $\rm MAE_{s}$$\downarrow$ $\rm OBO_{s}$$\uparrow$ $\rm MAE_{m}$$\downarrow$ $\rm OBO_{m}$$\uparrow$ $\rm MAE_{l}$$\downarrow$ $\rm OBO_{l}$$\uparrow$
-X3D [10] X3D 0.9105 0.1059 - - - - - -
-TANet [26] TANet 0.6624 0.0993 - - - - - -
-VideoSwinT [25] ViT 0.5756 0.1324 - - - - - -
-GTRM [15] I3D 0.5267 0.1589 - - - - - -
-RepNet [9] ResNet 0.5865 0.2450 0.7793 0.0930 0.5893 0.1591 0.4549 0.4062
-Zhang et al. [47] 3D-ResNext 0.8786 0.1554 - - - - - -
-TransRAC [14] ViT 0.4891 0.2781 0.5789 0.0233 0.4696 0.2955 0.4420 0.4375
-Li et al. [19] ViT 0.3841 0.3860 - - - - - -
-TadTR [24] $\dagger$ I3D 1.1314 0.0662 0.8364 0.0233 1.1591 0.0000 1.3106 0.1406
-ActionFormer [45] $\dagger$ I3D 0.4990 0.2781 0.4164 0.1628 0.3768 0.3409 0.6385 0.3125
-ReAct [34] $\dagger$ TSN 0.4592 0.3509 0.2805 0.1576 0.3037 0.4318 0.6862 0.3906
+ X3D [feichtenhofer2020x3d] X3D 0.9105 0.1059 - - - - - -
+TANet [liu2021tam] TANet 0.6624 0.0993 - - - - - -
+VideoSwinT [liu2022video] ViT 0.5756 0.1324 - - - - - -
+GTRM [huang2020improving] I3D 0.5267 0.1589 - - - - - -
+ RepNet [dwibedi2020counting] ResNet 0.5865 0.2450 0.7793 0.0930 0.5893 0.1591 0.4549 0.4062
+Zhang et al. [zhang2020context] 3D-ResNext 0.8786 0.1554 - - - - - -
+TransRAC [hu2022transrac] ViT 0.4891 0.2781 0.5789 0.0233 0.4696 0.2955 0.4420 0.4375
+Li et al. [li2024repetitive] ViT 0.3841 0.3860 - - - - - -
+ TadTR [liu2022end] $\dagger$ I3D 1.1314 0.0662 0.8364 0.0233 1.1591 0.0000 1.3106 0.1406
+ActionFormer [zhang2022actionformer] $\dagger$ I3D 0.4990 0.2781 0.4164 0.1628 0.3768 0.3409 0.6385 0.3125
+ReAct [shi2022react] $\dagger$ TSN 0.4592 0.3509 0.2805 0.1576 0.3037 0.4318 0.6862 0.3906
   Ours TSN 0.2809 0.4570 0.2411 0.1628 0.1792 0.5455 0.3776 0.5938
@@
 Ours ViT 0.2622 0.5430 0.2257 0.2558 0.2002 0.5909 0.3294 0.7031
-UCFRep dataset [47] is a subset of the UCF101 dataset [36], including fitness videos and daily life videos.
-Following previous work [14, 19], we do not use the train set but directly test our model on the test set to evaluate the model generalization ability.
+UCFRep dataset [zhang2020context] is a subset of the UCF101 dataset [soomro2012ucf101], including fitness videos and daily life videos.
+Following previous work [hu2022transrac, li2024repetitive], we do not use the train set but directly test our model on the test set to evaluate the model generalization ability.
 Metrics.
-Following previous works [14, 9, 47, 19], we compute two commonly used metrics, OBO and MAE, to evaluate the model performance.
-OBO (Off-By-One count error) measures the percentage that the predicted count is within the GT count ±$\pm 1$ range.
+Following previous works [hu2022transrac, dwibedi2020counting, zhang2020context, li2024repetitive], we compute two commonly used metrics, OBO and MAE, to evaluate the model performance.
+OBO (Off-By-One count error) measures the percentage that the predicted count is within the GT count $\pm 1$ range.
 MAE (Mean Absolute Error) measures the normalized absolute difference between the predicted and GT counts.
@@
 4.2 Implementation Details
-We implement our approach with different backbone video feature extractors respectively, including TSN [41], I3D [4], and ViT [8].
+We implement our approach with different backbone video feature extractors respectively, including TSN [wang2016temporal], I3D [carreira2017quo], and ViT [dosovitskiy2021an].
 For encoder-decoder Transformer, we set ${L}^{\text{enc}}=2$ and ${L}^{\text{dec}}=4$, both with 8-head attention mechanisms.
@@
 The length of the video input is set to $T=512$ frames without down-sampling.
-We utilize the AdamW optimizer [27] with a learning rate of $0.002$, a batch size of $64$, and train the model for $80$ epochs.
+We utilize the AdamW optimizer [loshchilov2017decoupled] with a learning rate of $0.002$, a batch size of $64$, and train the model for $80$ epochs.
 We set $\lambda_{\text{Hungarian}}=1.0$, $\lambda_{\text{L1}}=5.0$, $\lambda_{\text{gIoU}}=0.4$, $\lambda_{\text{ctrs}}=1.0$, confidence threshold $\alpha=0.2$.
@@
 4.3 Comparison to State-of-the-arts
-Table 2: Generalization comparison with SOTA TRC methods on UCFRep [47] dataset. MAE and OBO metrics for short-, medium-, and long-period actions are also reported.
+Table 2: Generalization comparison with SOTA TRC methods on UCFRep [zhang2020context] dataset. MAE and OBO metrics for short-, medium-, and long-period actions are also reported.
   Backbone $\rm MAE$$\downarrow$ $\rm OBO$$\uparrow$ $\rm MAE_{s}$$\downarrow$ $\rm OBO_{s}$$\uparrow$ $\rm MAE_{m}$$\downarrow$ $\rm OBO_{m}$$\uparrow$ $\rm MAE_{l}$$\downarrow$ $\rm OBO_{l}$$\uparrow$
-RepNet [9] ResNet 0.5336 0.2984 0.6219 0.1739 0.4825 0.3600 0.4996 0.5000
-TransRAC [14] ViT 0.6180 0.3143 0.6296 0.1951 0.5842 0.4250 0.6784 0.4118
-Li et al. [19] 3D-ResNext 0.5227 0.3500 - - - - - -
+ RepNet [dwibedi2020counting] ResNet 0.5336 0.2984 0.6219 0.1739 0.4825 0.3600 0.4996 0.5000
+TransRAC [hu2022transrac] ViT 0.6180 0.3143 0.6296 0.1951 0.5842 0.4250 0.6784 0.4118
+Li et al. [li2024repetitive] 3D-ResNext 0.5227 0.3500 - - - - - -
   Ours TSN 0.6016 0.2959 0.7069 0.0488 0.5777 0.4250 0.4039 0.5882
@@
 Ours ViT 0.5435 0.4184 0.5657 0.1951 0.4625 0.5500 0.6804 0.6471
-Table 3: Comparison of computational complexity. We benchmark all the methods with inference on the RepCountA dataset.
-       $T$        Params (M)        FLOPs (G)
-       RepNet [9]        64        20.47        92.40
-       TransRAC [14]        64        14.43        100.19
-       Ours        64        18.87        1.44
-       Ours        512        18.87        6.06
+Table 3: Comparison of computational complexity and inference time across different models.
+We evaluate the complexity and efficiency on varying frame lengths ($T$).
+OOM denotes “Out-of-Memory”, indicating that the model ran out of memory on a 32G GPU, and therefore, no inference time could be reported.
+    Model    $T$    Params (M)    FLOPs (G)    Inference Time (s)
+    RepNet [dwibedi2020counting]    64    43.98    163.44    3.62
+   512    48.94    6832.50    OOM
+   TransRAC [hu2022transrac]    64    42.28    582.02    2.43
+   512    48.22    8555.78    OOM
+        64    42.38    72.48    0.15
+   Ours    512    42.38    574.38    0.78
 Results on RepCountA dataset.
-We compare our proposed approach to the state-of-the-art methods on the RepCountA [14] dataset following previous work [14, 19] in Tab.
-1.
-We compare with the SOTA action recognition [10, 26, 25], action segmentation [15] methods (top block), and TRC [9, 47, 14, 19] approaches (second block).
-We further adapt recent DETR-style action detection approaches [24, 45, 34] for the TRC task (third block).
+We compare our proposed approach to the state-of-the-art methods on the RepCountA [hu2022transrac] dataset following previous work [hu2022transrac, li2024repetitive] in Tab.˜1.
+We compare with the SOTA action recognition [feichtenhofer2020x3d, liu2021tam, liu2022video], action segmentation [huang2020improving] methods (top block), and TRC [dwibedi2020counting, zhang2020context, hu2022transrac, li2024repetitive] approaches (second block).
+We further adapt recent DETR-style action detection approaches [liu2022end, zhang2022actionformer, shi2022react] for the TRC task (third block).
 We change their output layers accordingly and train them on RepCountA.
-As shown in Tab.
-1, our approach significantly outperforms the state-of-the-art methods across actions of varying lengths.
-Specifically, the similarity-matrix-based methods [9, 14, 19] suffer from quadratic computation complexity.
+As shown in Tab.˜1, our approach significantly outperforms the state-of-the-art methods across actions of varying lengths.
+Specifically, the similarity-matrix-based methods [dwibedi2020counting, hu2022transrac, li2024repetitive] suffer from quadratic computation complexity.
 To manage this, they employ a sparse sampling strategy, ensuring reasonable content coverage within a limited temporal context window.
 However, this approach results in inferior performance for short, rapid action instances, as it tends to overlook the cycles.
-Conversely, the action detection approaches [24, 45, 34] are primarily developed for detecting action instances specific to particular classes.
+Conversely, the action detection approaches [liu2022end, zhang2022actionformer, shi2022react] are primarily developed for detecting action instances specific to particular classes.
 To adapt them to the TRC task, we modify their output layer to incorporate class-agnostic supervision accordingly.
@@
 We also evaluate the generalization ability of our method.
-Following previous work [14, 19], we evaluate the model trained on the RepCountA dataset [14] on UCFRep [47] test set.
-Tab.
-2 shows that our approach generally outperforms existing works, with more significant improvements observed in longer-period actions.
+Following previous work [hu2022transrac, li2024repetitive], we evaluate the model trained on the RepCountA dataset [hu2022transrac] on UCFRep [zhang2020context] test set.
+Tab.˜2 shows that our approach generally outperforms existing works, with more significant improvements observed in longer-period actions.
 The results demonstrate the effectiveness and generalization capability of our method.
@@
 TransRAC represents the results by density map, and the final count value is obtained by summing the values in the density map.
-Figure 4: Visualization of failure case on RepCountA [14] dataset.
+Figure 4: Visualization of failure case on RepCountA [hu2022transrac] dataset.
 Due to the excessive zooming in, the legs of the human body are truncated, making a large difference in the action motion feature, and resulting in several missed cycle counts.
@@
 TransRAC represents the results by density map, and the final count value is obtained by summing the values in the density map.
-The vertical lines in the labels represent the time points at which the actions begin since only the starting point annotations are provided in UCFRep [47].
-Table 4: Effect of DAQ and ICL modules on RepCountA [14] dataset. (a) ablates DAQ strategy. (b) ablates the ICL strategy. (c) is our full model.
+The vertical lines in the labels represent the time points at which the actions begin since only the starting point annotations are provided in UCFRep [zhang2020context].
+Table 4: Effect of DAQ and ICL modules on RepCountA [hu2022transrac] dataset. (a) ablates DAQ strategy. (b) ablates the ICL strategy. (c) is our full model.
   $\rm MAE$$\downarrow$ $\rm OBO$$\uparrow$ $\rm MAE_{s}$$\downarrow$ $\rm OBO_{s}$$\uparrow$ $\rm MAE_{m}$$\downarrow$ $\rm OBO_{m}$$\uparrow$ $\rm MAE_{l}$$\downarrow$ $\rm OBO_{l}$$\uparrow$
@@
 We benchmark all the methods with inference on the RepCountA dataset.
-For fairness of comparison, we calculate the Floating Point Operations (FLOPs) in the counting module for all the methods, excluding the video feature extraction process.
-We set the batch size to be $1$ and clip length to be $64$ for each method.
-Tab.
-3 shows that our query-based method significantly reduces the computation load compared to the previous similarity-matric-based methods [9, 14], while also delivering superior accuracy.
-Importantly, the difference in computational demand is expected to widen with longer input clip lengths.
-This is because the computational complexity for similarity-matrix-based methods increases quadratically, whereas it grows linearly in our approach.
-Unfortunately, due to computational limitations, we are unable to obtain the results of RepNet [9] and TransRAC [14] when $T=512$.
+As shown in Tab.˜3, we present a detailed comparison including parameter count, FLOPs, and inference time, all measured on a single NVIDIA V100 32G GPU with a bacth size to be 1.
+Each method’s performance across two varying frame lengths ($T=64$ and $T=512$) is evaluated on the full model, to provide a more comprehensive assessment.
+Notably, prior methods like RepNet [dwibedi2020counting] and TransRAC [hu2022transrac] were originally designed to accept video inputs primarily at $T=64$ frames.
+When we attempted to process $T=512$ frames with their implementations, their computational complexity increased drastically.
+This is because both methods rely on a temporal similarity correlation matrix, causing their complexity to grow quadratically with the input length $T$.
+Consequently, they encountered “Out-of-Memory (OOM)” errors during execution, failing to obtain their inference time.
+In contrast, our method is significantly more lightweight and achieves demonstrably faster inference times compared to prior approaches.
+More importantly, its computational complexity scales linearly with increasing input video length ($T$), which robustly ensures the algorithm’s efficiency and allows it to maintain a very fast processing speed even when processing $T=512$ frames, as evidenced in Tab.˜3.
+Please refer to Supplementary for a decomposed model component complexity analysis.
 4.4 Qualitative Results
-We visualize the predictions of our approach and baseline method TransRAC [14] on the RepCountA [14] dataset in Fig.
-3.
-TransRAC [14] represents action cycles using density maps, and the final count value is obtained by summing the values in the density map.
+We visualize the predictions of our approach and baseline method TransRAC [hu2022transrac] on the RepCountA [hu2022transrac] dataset in Fig.˜3.
+TransRAC [hu2022transrac] represents action cycles using density maps, and the final count value is obtained by summing the values in the density map.
 However, this approach suffers from a lack of interpretability and finally results in miscounting.
 In contrast, our approach not only produces an accurate final count but also correctly localizes the action start and end positions (colored blocks) in most cases.
-In addition, our method exhibits robustness to changes in viewpoint, background noise, and sudden interruptions, e.g., in the second case of Fig.
-3, we accurately estimate the time positions even with viewpoint changes as well as a sudden interruption in the middle of the timeline.
+In addition, our method exhibits robustness to changes in viewpoint, background noise, and sudden interruptions, e.g., in the second case of Fig.˜3, we accurately estimate the time positions even with viewpoint changes as well as a sudden interruption in the middle of the timeline.
 Please refer to our supplementary video for more qualitative results.
-We present the video result of the first case in Fig.
-3, which shows that while the subject performs pull-ups, there are moments of talking and arm waving.
+We present the video result of the first case in Fig.˜3, which shows that while the subject performs pull-ups, there are moments of talking and arm waving.
 Our method selectively counts only the pull-up actions, as they are the primary focus and consistent with the main action instances of the video, excluding dissimilar actions like “waving”, validating the robustness and generalization ability of our method.
 In addition, our query-based representation offers excellent interpretability, making it easy to identify the issues in case of failures.
-For example, Fig.
-4 shows a typical failure case.
+For example, Fig.˜4 shows a typical failure case.
 Due to the excessive zooming in, the legs of the human body are truncated, making a large difference in the action motion feature, and resulting in several missed cycle counts.
-We further illustrate the generalization performance of the proposed method on the unseen UCFRep [47] test set in Fig.
-5.
+We further illustrate the generalization performance of the proposed method on the unseen UCFRep [zhang2020context] test set in Fig.˜5.
 We directly apply our trained model and do not use UCFRep training data.
@@
 We implement two ablated models to study the efficacy of the proposed DAQ and ICL designs.
-Tab.
-4 presents the results on RepCountA [14] using the TSN backbone.
+Tab.˜4 presents the results on RepCountA [hu2022transrac] using the TSN backbone.
 In ablation (a), we substitute the proposed DAQ module with a static action query, where the action queries $\mathbf{Q}^{\text{act}}$ are also learnable variables, instead of using the action queries filtered out by the encoder and query selection modules.
@@
 Confidence threshold $\alpha$.
-We report the impact of different confidence thresholds for classifying “repetitive actions” on the final counts in Fig.
-6.
-We illustrate the MAE (left) and OBO (right) curves for our method (green curve) and the TransRAC [14] approach (blue curve) across various confidence thresholds $\alpha$.
+We report the impact of different confidence thresholds, which range from 0 to 1 for classifying “repetitive actions” on the final counts in Fig.˜6.
+During inference, queries with prediction confidence scores below the threshold are classified as “others”, while the rest are considered “repetitive actions”.
+We illustrate the MAE (left) and OBO (right) curves for our method (green curve) and the TransRAC [hu2022transrac] approach (blue curve) across various confidence thresholds $\alpha$.
 The metrics for TransRAC are derived by binarizing its output density map and summing the results to obtain the final count.
-The results indicate that setting the threshold within the range of $0.2$ to $0.4$ yields similar performance for our method.
-Our method consistently outperforms TransRAC [14] (blue curve) by a considerable margin.
-Figure 6: Results of different confidence thresholds of our method and TransRAC [14].
+As the figure illustrates, our model performance drops significantly when the threshold exceeds 0.5, e.g.
+OBO becomes nearly zero, since most queries are mistakenly filtered out.
+This aligns with our expectations.
+Conversely, when the threshold is set between 0.3 and 0.5, the model yields consistently strong performance.
+However, setting the threshold too low also leads to performance degradation, as nearly all queries are then classified as “repetitive actions”, including irrelevant or noisy ones.
+This highlights the importance of selecting an appropriate threshold range to achieve optimal performance, as the results indicate that setting the threshold within the range of $0.2$ to $0.4$ yields consistently strong performance for our method.
+Besides, our method consistently outperforms TransRAC [hu2022transrac] (blue curve) by a considerable margin.
+Figure 6: Results of different confidence thresholds of our method and TransRAC [hu2022transrac].
 We depict the MAE (left) and OBO (right) curves of our method (green curve) and the TransRAC (blue curve) approach concerning different confidence thresholds.
@@
 Data Availability Statements.
-The RepCountA [14] and UCFRep [47] datasets that support the findings of this study are publicly available on GitHub at https://github.com/SvipRepetitionCounting/TransRAC and https://github.com/Xiaodomgdomg/Deep-Temporal-Repetition-Counting, respectively.
+The RepCountA [hu2022transrac] and UCFRep [zhang2020context] datasets that support the findings of this study are publicly available on GitHub at https://github.com/SvipRepetitionCounting/TransRAC and https://github.com/Xiaodomgdomg/Deep-Temporal-Repetition-Counting, respectively.